### v1

# Locret: Enhancing Eviction in Long-Context LLM Inference with Trained Retaining Heads

Abstract
Large language models (LLMs) have shown remarkable advances in supporting long-context comprehension and processing tasks. However, scaling the generation inference of LLMs to such long contexts incurs significant additional computation load, and demands a substantial GPU memory footprint to maintain the key-value (KV) cache of transformer-based LLMs. Existing KV cache compression methods, such as quantization, face memory bottlenecks as context length increases, while static-sized caches, such as selective eviction, suffer from inefficient policies. These limitations restrict deployment on consumer-grade devices like a single Nvidia 4090 GPU. To overcome this, we propose Locret, an efficient framework for long-context LLM inference that introduces retaining heads to evaluate the causal importance of KV cache units, allowing for more accurate eviction within a fixed cache size. Locret is fine-tuned on top of the frozen backbone LLM using a minimal amount of data from standard long-context SFT datasets. During inference, we evict low-importance cache units along with a chunked prefill pattern, significantly reducing peak GPU memory usage. We conduct an extensive empirical study to evaluate Locret, where the experimental results show that Locret outperforms the recent popular and competitive approaches, including InfLLM, Quantization, SirLLM, and MInference, in terms of memory efficiency and the quality of generated contents — Locret achieves over a 20×20$\times$20 × and 8×8$\times$8 × KV cache compression ratio compared to the full KV cache for Phi-3-mini-128K and Llama-3.1-8B-instruct. Additionally, Locret can be combined with other efficient inference methods, such as quantization and token merging. To the best of our knowledge, Locret is the first framework capable of deploying Llama-3.1-8B or similar models on a single Nvidia 4090 GPU, enabling 128K long-context inference without compromising generation quality, and requiring little additional system optimizations. The experimental code and the tool kit of Locret is provided at https://github.com/huangyuxiang03/Locret
1 Introduction
Large language models (LLMs) have revolutionized AI development and deployment (Zhao et al., 2023; Minaee et al., 2024). Recent advancements in LLMs’ ability to handle long-context tasks have further unlocked the potential of generative AI. State-of-the-art LLMs now support significantly extended context lengths, with GPT-4o (OpenAI, 2024) and Llama-3.1 (Dubey et al., 2024) handling 128K tokens, Yi (Young et al., 2024) and Claude-3 (Anthropic, 2024) supporting 200K tokens, and Gemini-1.5 (Reid et al., 2024) reaching 10 million tokens. These advances enable LLMs to tackle complex tasks like multi-hop reasoning (Li et al., 2024; Schnitzler et al., 2024), solving needle-in-a-haystack problems (Guerreiro et al., 2023; Wang et al., 2024a), and powering advanced LLM agents (Qin et al., 2023; Wang et al., 2024b) and AI-driven operating systems (Mei et al., 2024). However, deploying generative inference under long-context settings on consumer-grade GPUs requires innovative algorithmic and system optimizations to handle this new paradigm efficiently.
Compared to traditional short-context LLM inference, long-context LLM inference shifts the computing paradigm in two key ways: i) increased computational overhead for attention mechanisms: as context length grows, the computation required for obtaining attention scores increases quadratically, which results in a higher ratio of the computational budget in a transformer block; ii) higher memory footprint for key-value (KV) caching: longer contexts require larger KV caches, which dramatically increases the peak memory usage. These shifts demand innovative techniques to mitigate computational costs and manage memory usage effectively for long-context LLM inference.
Although various efforts have been made to overcome the bottleneck in LLM inference, these approaches fail to enable long-context inference on consumer-grade GPUs. Models with compact architectures, e.g. MiniCPM-128K (Hu et al., 2024b) and Phi-3-mini-128K (Abdin et al., 2024), reduce the computational load and memory usage, but cannot alleviate the KV cache burden in long-context scenarios. Similarly, techniques like LLM model weight quantization (Frantar et al., 2023; Lin et al., 2024; Ma et al., 2024), activation quantization (Dettmers et al., 2022; Xiao et al., 2023; Zhang et al., 2024c), or sparsification (Liu et al., 2023; Zhang et al., 2022) also fall short in effectively reducing the memory usage to a level supported by consumer-grade GPUs.
Recently, some specific optimizations have been proposed for long-context LLM inference. For example, sparse attention mechanisms (Jiang et al., 2024a; Ge et al., 2024; Lou et al., 2024) attempt to reduce runtime memory through conduct selected few calculation, and KV cache quantization (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024) reduces cache size by applying low-bit storage. These methods can only offer limited compression rate, of which the core issue is that the KV cache grows linearly with sequence length, and the above methods do not adequately address this problem. On the contrary, combining chunked prefill with token-dropping techniques (Xiao et al., 2024b; Yang et al., 2024) could offer a more effective solution, as it maintains a static-sized cache where the memory usage can be bounded. However, current token-dropping and cache eviction methods (Zhang et al., 2024e; Liu et al., 2024a; Yao et al., 2024), whose token importance manually designed according to the inner statistics, suffer from accuracy loss and performance degradation due to inaccuracies in estimating token importance — The weakening correlation between local and global importance as sequences grow exacerbates this issue. To address these limitations, we propose a lightweight training-based paradigm that provides more accurate token importance scoring to tackle the long-context LLM inference problem. We highlight our contributions below:
Contribution 1: We propose a lightweight training-based paradigm for selective KV cache eviction for long-context LLM inference, with an offline training cost $<$1 GPU hours. We tackle the problem of KV cache eviction by a learning based approach. We introduce the retaining heads, with a small number of additional parameters, fine-tuned on top of the frozen backbone LLM using a minimal amount of data from standard long-context SFT datasets to estimate the causal importance of each cache unit. Such a training paradigm is able to provide accurate token importance scoring prediction and can be integrated with other efficient inference algorithms, e.g., quantization and token merging.
Contribution 2: We provide an efficient inference system implementation for Locret. We integrate the retaining head mechanism into a chunked prefill inference framework, where we maintain a static-size cache set through evicting cache units with low predicted importance to limit the GPU memory usage. Locret is able to preserve the most important cache units with the trained retaining heads, enabling precise attention approximation without compromising the inference latency. Locret is also applicable to all transformer-based LLMs and various hardware, as it requires minimal modifications to the model’s inference process and only utilizes dense operators.
Contribution 3: We conduct an extensive evaluation of Locret, which illustrates that Locret can not only obtain a comparable performance but also maintain inference efficiency. Locret achieves over a 20×20$\times$20 × and 8×8$\times$8 × KV cache compression ratio for Phi-3-mini-128K and Llama-3.1-8B-instruct, enabling full comprehension of long contexts on consumer-grade devices. To the best of our knowledge, Locret is the first framework capable of deploying Llama-3.1-8B or similar models on a single Nvidia 4090 GPU, enabling 128K long-context inference without compromising generation quality, and requiring little extra system optimizations.
2 Related Work
Efforts in long-context LLM inference can be categorized by algorithm and system optimizations:
Algorithm optimizations. Optimizations aimed at reducing the size of the KV cache can generally be classified into three categories: quantization-based methods, token dropping methods, and sparsity-based methods. Quantization-based methods (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024; Zhang et al., 2024a), which store the KV cache in low-bit representations, require hardware support for these formats and may slow down inference due to the overhead of dequantization. Token dropping methods typically follow two main strategies: eviction or the use of an attention pool. Eviction-based approaches, such as H2(Zhang et al., 2024e), ScissorHands (Liu et al., 2024a), and SirLLM (Yao et al., 2024), rank tokens by certain statistical metrics to identify the most influential ones, discarding others to reduce memory usage. Attention pool-based methods (Nawrot et al., 2024; Rajput et al., 2024), such as StreamingLLM (Xiao et al., 2024b) and LoCoCo (Cai et al., 2024), compress multiple adjacent KV cache units into a single unit using a specially designed transformation. Sparsity-based methods (Ge et al., 2024; Jiang et al., 2024a; Yang et al., 2024; Lou et al., 2024; Lv et al., 2024) focus on leveraging the sparsity patterns of attention heads to reduce both computation and I/O. The combination of these approaches can be further enhanced by identifying specific attention patterns for each head and layer (Ge et al., 2024). For surveys of these methods, please refer to (Yuan et al., 2024; Kang et al., 2024; Shi et al., 2024).
System optimizations. The challenge of long-context inference can also be alleviated from a system-level perspective. Offloading-based methods (Sheng et al., 2023; Xiao et al., 2024a; Wu et al., 2024; Lee et al., 2024) store the KV cache in CPU memory, retrieving only the most relevant chunks to the GPU before computing a new chunk. This approach reduces peak GPU memory usage, though at the cost of slower inference. Hardware-aware algorithms, such as flash attention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) and page attention (Kwon et al., 2023), exploit GPU architecture (Ghorpade et al., 2012) to enable more efficient runtime memory management. In addition, reimplementing inference infra-structure in a more efficient programming language (llama.cpp, ; llama2.c, ; rustformers, ), or adopting disaggregated inference (Jiang et al., 2024b; Zhong et al., 2024; Qin et al., 2024; Hu et al., 2024a), can greatly enhance inference efficiency. Algorithmic optimizations can be seamlessly integrated into such systems (Agrawal et al., 2023; Lee et al., 2024). For instance, KTransformers (KVCache.AI, 2024) adopts the chunked offloading technique from InfLLM (Xiao et al., 2024a). However, system optimizations primarily focus on extending context length by leveraging hardware resources, rather than directly reducing the size of the KV cache.
3 Locret
3.1 Preliminaries
Transformer architecture. We define the model inference of transformer-based LLMs as follows. Given a token sequence $t_{1},t_{2},\cdots,t_{n}$, we denote the output hidden state of layer $i$ as $\textbf{H}^{(i)}$ and $\textbf{H}^{(0)}$ is the embeddings. Each transformer layer consists of an attention layer and an MLP layer. We assume the model follows a grouped-query attention (GQA) architecture (Ainslie et al., 2023), with $h$ query heads and a group size of $g$. For multi-head attention (MHA), $g$ is set to 1. The attention score of layer $i$’s head $j$ is calculated by $\mathbf{A}_{j}^{(i)}=\text{softmax}\left(\mathbf{Q}_{j}^{(i)}{\mathbf{K}_{\lceil{j/g}\rceil}^{(i)T}}/\sqrt{d_{m}}\right)\cdot\mathbf{V}_{\lceil{j/g}\rceil}^{(i)}$, where $d_{m}$ represents the hidden size for each head and $\left[\mathbf{Q}_{j}^{(i)},\mathbf{K}_{j}^{(i)},\mathbf{V}_{j}^{(i)}\right]=\mathbf{H}^{(i-1)}\cdot\left[{\mathbf{W}_{j}^{Q(i)}},{\mathbf{W}_{j}^{K(i)}},{\mathbf{W}_{j}^{V(i)}}\right]$. Next, we compute $\mathbf{A}^{(i)}=\left[\mathbf{A}_{1}^{(i)},\cdots,\mathbf{A}_{h}^{(i)}\right]\cdot\mathbf{W}^{O(i)}$, finally followed by $\mathbf{H}^{(i)}=\text{MLP}(\mathbf{A}^{(i)})$.
KV cache and chunked prefill. During the prefill stage, all prompt tokens are processed in a single forward pass, where $\mathbf{Q}^{(i)}$, $\mathbf{K}^{(i)}$, and $\mathbf{V}^{(i)}$ each have a sequence length of $n$. In the decoding stage, only a single token is processed across layers, utilizing the KV cache units to reduce computation. Chunked prefill is a method for reducing peak memory consumption by processing tokens in chunks over multiple passes, with the assistance of the KV cache. Taking both KV cache and chunked prefill into account, the attention calculation can be modified as Equation 1, where $B$ represents the number of tokens processed in a single model pass. For decoding, $B=1$, while for chunked prefill, $B$ corresponds to the chunk size. We denote the attention output for tokens $n+1,n+2,\cdots,n+B$ as $\mathbf{A}[n+1\colon n+B]$, and the attention output for the $k$-th token as $\mathbf{A}[k]$.
$\mathbf{A}{[n+1\colon n+B]}_{j}^{(i)}=\text{softmax}\left(\frac{\mathbf{Q}{[n+1\colon n+B]}_{j}^{(i)}{\mathbf{K}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)T}}}{\sqrt{d_{m}}}\right)\cdot\mathbf{V}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)}.$ (1)
Cache eviction. Cache eviction in long-context inference is defined as follows. Here, we slightly abuse the notation of heads and layers, and treat the key-value vector pair of a single token within one head as the smallest cache unit. We denote the cache unit for the $k$-th token as $c_{k}=(\mathbf{K}[k],\mathbf{V}[k])$. Assume a memory budget $b$, representing the maximum number of cache units that can be stored in GPU memory at any given time. The abstract form of attention can then be written as $c_{k}=f(c_{1},c_{2},\cdots,c_{k-1})$. With limited cache capacity, this calculation can only be approximated by $\tilde{c}_{k}=f(\tilde{c}_{p_{1}},\tilde{c}_{p_{2}},\cdots,\tilde{c}_{p_{b^{\prime}}})$, where $b^{\prime}\leq b$, and $p_{1},p_{2}\cdots,p_{b^{\prime}}\in\{1,2,\cdots,k-1\}$. Intuitively, the number of prior cache units involved cannot exceed the memory budget. When $b^{\prime}=b$, indicating the cache is full, one cache unit must be evicted. We select the unit to be evicted using a policy $p_{v}=\text{Policy}(\tilde{c}_{p_{1}},\cdots,\tilde{c}_{p_{b}};\tilde{c}_{k})$. In such stated problem, the key challenge of cache eviction is to develop an effective policy function that minimizes the approximation error $\|\tilde{c}_{k}-c_{k}\|$.
Figure 1: The framework of Locret. “$\mathbf{R}$” represents the retaining head. $\text{P}_{i}$ and $\text{A}_{i}$ correspond to the $i$-th prompt token and answer token. “t” represents the time step in chunked prefill, “$b$” represents the budget size, and “$n_{s}$” represents the length of the stabilizers.
3.2 Locret  Framework
Locret is a training-based KV cache compression framework that works in conjunction with chunked prefill. As illustrated in Figure 1, Locret operates in two stages: training and inference. In the training stage, we modify the original LLM by appending a retaining head $\mathbf{R}$ to each attention module. We then train the retaining heads $\mathbf{R}$ while keeping the LLM backbone frozen. During chunked prefill inference, the retaining heads $\mathbf{R}$ are used to calculate the importance of each cache unit in the chunk. We retain the cache units with higher scores, along with stabilizers (i.e., the last tokens), in the cache pool located in GPU memory. Through this process, the retaining heads $\mathbf{R}$ learn and predict the patterns discovered by existing methods, as detailed in Appendix G.
From a mathematical perspective, cache eviction is performed by assigning each cache unit an importance score that reflects its influence on the calculation of subsequent tokens. We refer to this estimation as the causal importance score (CIS) since it is computed in a causal manner. The CIS of cache unit $k$ is calculated as $s_{k}=S(c_{1},c_{2},\cdots,c_{k})$. By applying top-$b$ sparse attention based on the CIS, we can ensure that the trace (i.e. retaining and eviction) of each cache unit can be fit within a cache with a given memory budget. Further details can be found in Appendix F.
However, since not all tokens can be stored in the cache simultaneously, calculating the actual CIS on-chip is impractical. Instead, we use a heuristic approximation for CIS, defined as follows: $\tilde{s}_{k}=S(\tilde{c}_{p_{1}},\tilde{c}_{p_{2}},\cdots,\tilde{c}_{p_{b^{\prime}}})$, where $b$ is the cache budget, all $\tilde{c}_{p_{i}}$ are approximated cached units, and $b^{\prime}\leq b$. We hypothesize that if the scoring function for causal importance is sufficiently accurate, it will consistently select the most critical cache units, resulting in a negligible difference between the heuristic and the actual score. Thus, we use the terms heuristic CIS and actual CIS interchangeably.
3.3 Training the Retaining Heads
In this section, we introduce Locret’s model architecture modifications and the corresponding training recipe. We add additional parameters to compute the CIS $s_{k}$ (or $\tilde{s}_{k}$ for on-chip inference) with respect to all previous cache units. Specifically, we inject a retaining head, consisting of a small MLP, into each layer. The retaining head predicts the CIS for each head of the corresponding layer based on the concatenation of $[\mathbf{Q},\mathbf{K},\mathbf{V}]$. Formally, with a slight abuse of notation, let the retaining head for layer $i$ be denoted as $\mathbf{R}$. The CIS at head $j$ of layer $i$ is then calculated as: $\mathbf{\tilde{S}}=\mathbf{R}([\mathbf{Q},\mathbf{K},\mathbf{V}])=\sigma([\mathbf{Q},\mathbf{K},\mathbf{V}]\mathbf{W}_{1})\mathbf{W}_{2}$. In this equation, $\mathbf{W}_{1}\in\mathbb{R}^{(d_{m}+2d_{kv})\times d_{\mathbf{R}}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{\mathbf{R}}\times\frac{h}{g}}$ are the tunable parameters of $\mathbf{R}$, $\sigma$ is the activation function and $\mathbf{\tilde{S}}[k]_{j}$ is the predicted CIS of the $k$-th token at head $j$ of layer $i$. This architecture implies that the importance estimation for a single head is not performed in isolation but rather considers all heads together. Note that for GQA models, there are only $h/g$ output values corresponding to the number of heads in the KV cache.
We train the retaining head $\mathbf{R}$s on a small Question-Answer (QA) supervised fine-tuning (SFT) dataset, where each entry consists of a single prompt and one answer. We define the CIS $s_{k}$ for the $k$-th token as the maximum attention score, before softmax, from all the answer tokens toward the $k$-th token. Formally, for the $k$-th token at head $j$ of layer $i$, we approximate the predicted value $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ to the ground truth $\mathbf{S}[k]_{j}^{(i)}:=\max_{p}\left(\mathbf{Q}_{j}^{(i)}\mathbf{K}_{j}^{(i)T}\right)_{p,k}$, where $n_{q}(d)\leq p\leq n_{q}(d)+n_{a}(d)$, and $n_{q}(d)$ and $n_{a}(d)$ represent the lengths of the prompt and answer in data $d$, respectively. For an MHA model with $L$ layers and $h$ heads, the training objective is described in Equation 2. For GQA models, we take the maximum attention score before softmax across different query heads within the same group as the ground truth for the corresponding KV head.
$\operatorname*{arg\,min}_{\mathbf{W}_{1}^{(i)},\mathbf{W}_{2}^{(i)},i=1,2\cdots,L}\mathbb{E}_{d\in\mathcal{D}}\left[\sum_{i=1}^{L}\sum_{j=1}^{h}\sum_{k=1}^{n_{q}(d)}\mathcal{L}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)\right]$ (2)
The training loss consists of a regression loss and a smoothing loss. We apply the Smooth-$\mathcal{L}_{1}$ norm between the predicted values and the ground truth. Since important segments in natural language typically consist of adjacent tokens, we also apply the $\mathcal{L}_{2}$ norm between each pair of adjacent predicted values to enforce smoothness. The complete training loss for Locret is given by Equation 3.
$\mathcal{L}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)=\text{Smooth-}\mathcal{L}_{1}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)+\alpha\mathcal{L}_{2}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{\tilde{S}}[k+1]_{j}^{(i)}\right)$ (3)
Figure 2: L-Eval scores with different intermediate size of retaining head $d_{\mathbf{R}}$.
Table 1: L-Eval scores of Locret trained on various datasets.
Dataset LongAlpaca LongAlign Anti-Haystack
L-Eval 21.33 22.00 20.72
From our observations, the training of Locret exhibits strong robustness. The performance variations shown in Figure 2 and Table 2 are minimal, despite changes in $d_{\mathbf{R}}$ and the dataset. Details can be found in Appendix D. Training statistics, including loss dynamics, are recorded in Appendix H.
3.4 Inference Implementation of Locret
Input: Model $\mathbf{M}$, Prompt tokens $x$, Local length $n_{loc}$, Stablizer length $n_{s}$, Budget $b$, Chunk size $B$
Output: Generated tokens $x_{gen}$
// Leave the last $n_{loc}$ out to make sure they are not evicted.
chunk_positions $\leftarrow$ split_chunk(00, $x$.length() $-n_{loc}$, $B$)
K_cache, V_cache, score_cache $\leftarrow$ [], [], []
for chunk $\in$ chunk_positions do
      begin_pos, end_pos $\leftarrow$ chunk.begin_pos, chunk.end_pos
[…]
      K_cache $\leftarrow$ Concat(K_cache, K_chunk)
      V_cache $\leftarrow$ Concat(V_cache, V_chunk)
      score_cache $\leftarrow$ Concat(score_cache, score_chunk)
      if chunk is not the last chunk then
           // Keep the last $n_{s}$ caches to maintain higher context continuity.
           score_cache[score_cache.length()-$n_{s}$:score_cache.length()] $\leftarrow$+$\infty$ + ∞
           end if
          indices $\leftarrow$ top-$b$(score_cache).indices
           K_cache, V_cache, score_cache = K_cache[indices], V_cache[indices], score_cache[indices]
           end for
          K_cache, V_cache, score_cache $\leftarrow\mathbf{M}$($x$[$x$.length()$-n_{loc}$:$x$.length()], K_cache, V_cache)
           $x_{gen}\leftarrow\mathbf{M}$.generate(K_cache, V_cache)
           return $x_{gen}$
Algorithm 1 Locret Inference
During the inference stage, we use the chunked prefill pattern and perform cache eviction based on the predicted CIS. Since the predicted value $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ depends solely on $\mathbf{Q}[k]_{j}^{(i)}$, $\mathbf{K}[k]_{j}^{(i)}$, and $\mathbf{V}[k]_{j}^{(i)}$, and because attention in decoder-only models is causal, $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ remains consistent once calculated. Thus, we store the KV cache units along with their corresponding causal importance values. When the cache is full, we evict the units with lower causal importance values, as they are deemed less useful for future computations. Cache eviction introduces context discontinuity, meaning some cache units at certain positions may be absent. This can degrade generation quality and increase the error between the predicted and accurate CIS, as LLMs are typically not trained on such contexts. To mitigate this, we retain the last tokens of the current chunk at each step of the chunked prefill process, ensuring a local and continuous context to minimize errors. To demonstrate the effectiveness of this design, we perform an ablation study on the length of stabilizers $n_{s}$, shown in Figure 3. Smaller $n_{s}$ results in severe performance degradation, and the model fails entirely when stabilizers are absent, as context discontinuity leads to instability in CIS prediction, causing errors in cache eviction and amplifying errors in hidden states. More details are discussed in Appendix E.
Figure 3: R.Number with different stabilizer lengths $n_{s}$. (a) Task accuracy under different $n_{s}$. (b) Maximum absolute error of the last hidden state. (c) Mean absolute error of the predicted CIS. We conduct this experiment on entries 101-120 of R.Number using the Phi-3-mini-128K backbone.
We maintain a cache pool with a capacity of $b$ cache units, discarding units that exceed this limit. For each chunk, the model processes the chunked input tokens alongside the current cache pool. The newly generated KV pairs and their predicted scores are then concatenated with the existing cache. Once the cache pool is full, only the $b$ cache units with the highest CIS values are retained. At each chunked prefill step, except for the final step, we retain the stabilizers, i.e. the last $n_{s}$ cache units. Additionally, we do not compress the last $n_{loc}$ tokens of the prompt, as they are critical for maintaining high generation quality due to their strong correlation with the query. Finally, the answer is generated according to the compressed KV cache. Algorithm 1 provides the pseudocode for Locret, where we formally describe the Locret inference process.
The GPU memory usage during Locret inference can be effectively bounded. GPU memory for KV cache storage is limited to $\mathcal{O}(b+n_{loc})$, and the runtime memory usage of the attention mechanism is bounded by $\mathcal{O}(B\times(b+B+n_{loc}))$. For comparison, while processing an input with $n$ tokens, full attention prefill requires $\mathcal{O}(n)$ for KV cache storage and $\mathcal{O}(n^{2})$ for runtime memory, whereas chunked prefill requires $\mathcal{O}(n)$ for KV cache storage and $\mathcal{O}(nB)$ for runtime memory consumption.
4 Experiments
In this section, we present the experiments conducted to evaluate the proposed framework, Locret, aiming to address the following questions:
(Q1) Can Locret obtain better end-to-end task performance compared to popular and competitive long-conetext inference approaches within similar or less peak memory?
(Q2) Can Locret improve inference speed compared to other approaches?
(Q3) Can Locret be combined with other efficient inference methods?
4.1 Experimental Setup
Models and training dataset. We evaluate Locret on two long-context LLMs: Phi-3-mini-128K (Abdin et al., 2024) and Llama-3.1-8B-instruct (Dubey et al., 2024). Both models can process up to 128K context tokens and follow MHA and GQA architectures, respectively. Phi-3-mini-128K has 3.8B parameters, while Llama-3.1-8B-instruct contains 8B parameters, making both suitable for deployment on consumer-grade devices. We inject retaining heads $\mathbf{R}$ into each layer, setting the intermediate size $d_{\mathbf{R}}$ to 1024 for both models. The retaining heads are trained on the LongAlpaca dataset (Chen et al., 2024) for 3000 steps. Training Locret is lightweight, with the tunable parameters comprising 8% and 2.5% of the total for the two models, respectively. The complete training process takes 0.47 and 0.80 GPU hours on a single A800 GPU for each corresponding model. The cache budget $b$ is set to 6000 for Phi-3-mini-128K and 16384 for Llama-3.1-8B-instruct. More details on hyperparameters for both training and inference, as well as the system environment, can be found in Appendix A.
[…]
Baselines. As discussed in Section 2, existing algorithms for memory-efficient long-context inference can be categorized into offloading-based, quantization-based, token-dropping, and sparsity-based methods. For each category, we select one representative method as the baseline. We compare Locret against full attention inference (denoted as FullAttn), InfLLM (Xiao et al., 2024a), KV cache quantization (Turganbay, 2024), SirLLM (Yao et al., 2024), and MInference (Jiang et al., 2024a). For quantization, we use Hugging Face Quanto (Hugging-Face, ) implementation, referring to the 2-bit quantization method as HF-2bits. We omit HF-4bits and benchmark this combination in Section 4.4. We do not include attention pool-based token-dropping methods in this benchmark, as they are orthogonal to our approach; further discussion is provided in Section 4.4. Detailed introductions to the selected baselines can be found in Appendix A. We also discuss the comparison between the trained Locret and the randomly initialized retaining heads $\mathbf{R}$ in Appendix B.
4.2 End-to-end Benchmark
[…]
In the L-Eval tests, all methods show some degree of performance degradation. Nevertheless, Locret achieves the best overall result, obtaining the highest score on most tasks. L-Eval is a shorter but more complex dataset, where SirLLM performs particularly well. Quantization fails on most tasks, resulting in the lowest overall score. Both InfLLM and MInference suffer significant performance drops compared to FullAttn inference. Locret consistently surpasses all competitors.
[…]
Figure 4: Memory Statistics vs. Task Performance. The red lines correspond to the theoretical size of the model weights, while the blue lines represent the total theoretical size of the model weights and the full KV cache without any compression. The purple lines indicate the accuracies of FullAttn. “Total Memory” represents the total memory usage of both GPU and CPU memory.
[…]
Method R.PassKey R.Number E.Sum E.QA E.MC Z.QA E.Dia C.Debug M.Find Avg.$\uparrow$
[…]
FullAttn 98.64 97.12 17.92 11.16 55.46 14.83 8.00 23.10 17.43 38.18
InfLLM 100.00 97.12 14.35 4.97 38.86 11.04 3.50 25.38 15.14 34.48
HF-2bits 0.00 0.00 13.80 1.44 1.75 0.20 0.50 0.00 0.57 2.03
SirLLM 3.39 3.39 21.06 6.32 44.98 11.99 5.00 22.34 21.71 15.58
MInference 99.32 95.93 14.44 8.11 40.61 10.60 9.00 15.48 15.43 32.25
Locret 100.00 97.46 16.82 7.61 46.29 11.31 10.00 27.92 29.71 34.73
[…]
FullAttn 100.00 99.32 26.79 15.06 68.12 13.40 17.00 20.56 34.00 43.81
InfLLM 100.00 100.00 24.24 14.21 51.97 10.76 11.00 26.25 35.71 41.57
HF-2bits 36.78 6.95 8.77 4.05 27.95 3.09 5.50 13.20 22.00 14.25
SirLLM 1.69 1.69 25.60 8.95 55.46 10.38 9.50 23.10 3.71 15.56
MInference 100.00 98.47 20.64 14.35 59.83 12.20 20.50 25.89 35.43 43.03
Locret 100.00 99.49 27.28 20.90 58.82 11.85 13.00 27.16 32.86 43.48
Method CodeU NQ CUAD NarrativeQA QMSum SPACE Avg.$\uparrow$
Phi-3-mini-128K on L-Eval
FullAttn 8.89 59.14 30.34 17.59 16.05 14.51 24.42
InfLLM 5.56 34.32 14.53 14.80 13.31 14.81 16.22
HF-2bits 0.00 1.69 6.40 2.04 2.73 3.34 2.70
SirLLM 8.89 37.92 20.89 14.51 13.70 14.46 18.40
MInference 7.78 25.21 26.64 15.14 15.78 14.87 17.57
Locret 8.89 51.49 22.23 16.42 14.86 14.06 21.33
Llama-3.1-8B-instruct on L-Eval
FullAttn 10.0 66.84 38.91 23.11 18.76 16.86 29.08
InfLLM 6.67 54.77 33.76 20.35 17.62 16.73 24.98
HF-2bits 1.11 29.79 18.98 9.46 14.02 13.73 14.52
SirLLM 5.56 58.00 35.41 21.21 17.32 16.44 25.66
MInference 7.78 31.80 36.93 19.44 18.14 16.76 21.81
Locret 8.89 63.03 37.21 23.59 18.17 16.87 27.96
4.3 Speed Test on Real Consumer-Grade Devices
[…]
R.PassKey is a task where the model retrieves a 5-digit number from a large amount of irrelevant text, a task we believe to be relatively simple for humans. Thus, we consider the task to have failed if the accuracy falls below 95%. As shown in Table 3, aside from the settings that fail on this task, Locret achieves the highest inference speed among all methods that can correctly process R.PassKey. Due to its MHA architecture, Phi-3-mini-128K has a larger KV cache, which leads to failures for both HF-2bits and MInference. Storing the full KV cache on a single 4090 GPU is infeasible, as it requires 48GB of memory. Although the quantized KV cache is reduced to 6GB, the converting processes between representations requires significant GPU memory for intermediate states, resulting in the failure of HF-2bits. While InfLLM can run in memory-limited scenarios, its offloading process slows down inference, with I/O becoming the bottleneck in attention calculation. SirLLM fails due to its inaccurate eviction policy, which cannot correctly identify the 5-digit number. In the GQA model (Llama-3.1-8B-instruct), which has a smaller KV cache, the quantized cache can fit within the GPU memory. However, the quantization and dequantization processes become the bottleneck, leading to significantly slower speeds. The performance of InfLLM, SirLLM, and MInference is similar to that seen with Phi-3-mini-128K. Although MInference benefits from faster encoding speeds, it fails on this task because it cannot process the entire input sequence at once. Locret strikes a balance between inference speed and performance, making it a far more suitable solution for long-context scenarios on consumer-grade devices.
Table 3: Executing R.PassKey on an Nvidia 4090. “tok/s” represents the inference speed, “C.Len” stands for the context length after truncation, and “Acc.” represents task accuracy. “Success” shows whether a setting has $\geq$95accuracy. The highest score among 128K context is marked in bold.
Method FullAttn InfLLM HF-2bits SirLLM MInference Locret HF-2bits MInference
Phi-3- mini-128K tok/$\uparrow$ - 2276.38 - 2352.20 - 5080.85 1098.51 4099.92
C.Len.$\uparrow$ 128K 128K 128K 128K 128K 128K 30K 14K
Acc.$\uparrow$ OOM 99.83 OOM 1.69 OOM 100.00 0.00 13.56
Success
Llama-3.1- 8B-instruct tok/$\uparrow$ - 2287.66 1365.51 1589.75 - 3209.10 3680.06 5135.74
C.Len.$\uparrow$ 128K 128K 128K 128K 128K 128K 30K 25K
Acc.$\uparrow$ OOM 100.00 35.59 1.69 OOM 100.00 26.78 20.34
Success
4.4 Orthogonality to Other Methods
We evaluate the combination of Locret with quantization and token merging to address Q3.
Table 4: Quantization with FullAttn and Locret. “M” represents Method and “$-\Delta$” represents the gap of average L-Eval score.
Setting M M-4bits $-\Delta$
M=FullAttn 29.08 28.52 0.56
M=Locret 27.96 27.11 0.85
Table 5: The average L-Eval scores of LoCoCo, Locret, and the combination of LoCoCo and Locret.
Method LoCoCo Locret Combination
L-Eval 26.01 27.96 28.70
KV cache quantization. According to Zhang et al. (2024d), eviction-based methods like H2O struggle with compatibility when combined with KV cache quantization. Quantization introduces significant disturbance in the estimation of heavy-hitters, leading to severe performance degradation. However, Locret is not affected by such issues and can be combined with quantization while maintaining most of its performance. Here, we compare the performance degradation caused by quantization on Locret with that of the full attention method using the same metrics. We use Quanto as the quantization backend and report the average L-Eval score with Llama-3.1-8B-instruct as the model backbone. Table 4.4 shows that the performance drop caused by quantization on Locret is only slightly higher than that observed with the full attention method, indicating that Locret is a quantization-friendly approach. More details of the experiment are provided in Appendix C.1.
Token merging. As described in Section 2, token dropping can also be implemented through an attention pool. Attention pool-based methods (Xiao et al., 2024b; Cai et al., 2024; Mu et al., 2024; Munkhdalai et al., 2024) merge adjacent tokens or cache units into an attention pool, maintaining a static cache size. These methods are orthogonal to Locret , as the evicted tokens can be merged into a small cache pool and retained in GPU memory. We conduct the following experiment to demonstrate that Locret can serve as an effective plug-in scoring function within such frameworks, enhancing performance without increasing memory budget. We select LoCoCo (Cai et al., 2024) as a representative of the latest attention pool-based methods. LoCoCo maintains a cache set consisting of two parts: the heavy hitters and the convolved non-heavy hitters. During each chunked prefill step, LoCoCo first identifies a set of heavy hitters according to H2(Zhang et al., 2024e), then applies 1-D convolution to the non-heavy hitters to compress them into a static size. By replacing H2O’s heavy-hitter scoring function with Locret, we retain the cache units with high CIS and convolve the others. We compare this combination with standalone LoCoCo and Locret on L-Eval using the Llama-3.1-8B-instruct backbone and report the average score across all selected tasks. As shown in Table 4.4, Locret achieves a higher score than LoCoCo, and the combined algorithm outperforms both standalone methods. This suggests that Locret provides a more accurate scoring function compared to H2O, and the two methods complement each other, demonstrating their orthogonality. Further details of the experiment are provided in Appendix C.2.
5 Conclusion
We propose Locret, a lightweight training-based method that enables memory-efficient inference of long contexts on consumer-grade devices. Locret introduces a few additional parameters (retaining heads) to predict the CIS of each cache unit during chunked prefill and performs cache eviction based on the predicted CIS. We conduct extensive experiments across different models and multiple datasets to compare Locret with major efficient inference techniques, including offloading, quantization, eviction, and sparse attention. Results show that Locret outperforms all baselines, using less GPU memory and without requiring offloading to CPU memory. Furthermore, Locret can be combined with other efficient inference methods, such as KV cache quantization and attention pool-based token dropping. Locret maintains high inference throughput on consumer-grade devices and demonstrates robustness to variations in the intermediate size of retaining heads and training data. The complete framework of Locret, including both training and inference, highlights its suitability for low-resource computing scenarios. Locret can be applied to various application scenarios, such as end-side multi-modal model inference and context compression during disaggregated inference, which we plan to explore in future work. We also aim to test Locret on additional model architectures, including multi-latent models and models of varying sizes. Additionally, we are highly interested in integrating Locret into end-side deployment infra-structures to enhance inference efficiency, particularly for devices like smartphones and personal computers.
Acknowledgment
[…]
Appendix A Hyperparameters, Environment and Baselines
A.1 Training
During the training stage, we first insert retaining head $\mathbf{R}$s to each layar. A retaining head is a small FFN consist of two linear transformations, and the non-linear function is aligned with other non-linears of the conresponding model, with an intermediate size of 1024. We train the appended retaining head $\mathbf{R}$s on the LongAlpaca for 3000 steps with batch size set to 1 and maximum sequence length set to 10240. We use the AdamW scheduler (Loshchilov, 2019) and the learning rate is set to 5e-4. We conduct the training with a linear learning rate scheduler, whose warmup step number is set to 2000. The balance factor between two training loss $\alpha$ is set to 0.0025.
A.2 Inference
Table 6: Hyperparameters in Locret’s inference stage.
Model $b$ $B$ $n_{s}$ $n_{loc}$
Phi-3-mini-128K 6000 3072 2500 100
Llama-3.1-8B-instruct 16384 1024 2500 100
The inference hyperparameters of Locret is listed in Table 6. Here, we follow the notations in Algorithm 1. $b$ stands for the cache budget, $B$ is the chunk size of chunked prefill, $n_{s}$ is the length of stabilizers, and $n_{loc}$ represents the length of locally retained tokens at the end of the input sequence.
Hyperparameters of other baselines are as follows. For InfLLM, we use the recommended settings for Llama-3 to evaluate Llama-3.1. Since there is no recommendations of Phi-3-mini-128K, we use the settings for MiniCPM, whose architechture and size is similar to Phi-3-mini-128K, to conduct all the experiments. For Quantization, we use the official implementation (Quanto backend) of Hugging Face. For SirLLM, we set the start size to 4, recent size to 1000 for both models. We set the token entropy size to 6000 and 16384 for Phi-3-mini-128K and Llama-3.1-8B-instruct respectively. The chunk size of chunked prefill is also 3072 and 1024 for the corresponding model. For MInference, we utilize the recommended settings for both models.
A.3 System Environment
For all the experiments except the 4090 experiments in Section 4.3, we use a workstation with 8$\times$×Nvidia A800/H800 GPUs and 104 Intel(R) Xeon(R) Platinum 8470 CPUs. The device has 1.0 TB CPU memory. The operating system is Red Hat 4.8.5. We conduct all experiments except the full attention full KV cache inference on a single GPU, and 2 GPUs for full attention settings.
For Section 4.3, we conduct the experiments on a single Nvidia 4090 GPU. The device has 512 AMD EPYC 9754 128-Core Processors and 1.0 TB CPU memory. GPUs and CPUs are connected through PCIe Gen 4, which has 16GT/s transmission speed. The operating system is Ubuntu 9.4.0.
A.4 Baselines
We compare Locret with full attention inference, InfLLM, Quantization, SirLLM and MInference. FullAttn inference is performed using vllm (Kwon et al., 2023), which includes automatic tensor parallelism. InfLLM is a representative of the offloading-based methods, where the full KV cache is offloaded to CPU, and the most relavant blocks are retrieved to GPU during inference. For quantization method, we use the Hugging Face implementation of 2-bits KV cache quantization, which is inspired by Liu et al. (2024b), where quantization is conducted along channels instead of tokens. We denote this method as HF-2bits. SirLLM is an eviction-based token dropping algorithm, where tokens with low token-entropy is evicted once the cache is fullfilled. MInference is a typical method of reducing peak GPU memory consumption through rule-based sparse attention, but it does not reduce the size of KV cache. Note that InfLLM, HF-2bits and SirLLM does not have official implementation on Phi-3-mini-128K, thus we implement these three methods according to the original algorithm. We only use the short factor of RoPE for InfLLM, and no further model modification is conducted for HF-2bits and SirLLM.
Appendix B The Effect of Training
[…]
Method R.PassKey R.Number E.Sum E.QA E.MC Z.QA E.Dia C.Debug M.Find Avg.
Random 0.00 34.00 5.09 2.68 18.34 1.54 0.00 13.71 2.57 4.92
Locret 100.00 97.46 16.82 7.61 46.29 11.31 10.00 27.92 29.71 34.73
[…]
Appendix C Orthogonality to Other Methods
C.1 Combination with Quantization
Table 8: L-Eval scores of FullAttn, FullAttn-4bits, Locret and Locret-4bits. (Detailed)
Llama-3.1-8B-instruct on L-Eval
Method CodeU NQ CUAD NarrativeQA QMSum SPACE Avg.$\uparrow$
FullAttn 10.0 66.84 38.91 23.11 18.76 16.86 29.08
FullAttn-4bits 7.78 66.64 38.25 22.76 18.85 16.84 28.52
Locret 8.89 63.03 37.21 23.59 18.17 16.87 27.96
Locret-4bits 4.44 63.22 36.95 22.80 18.43 16.81 27.11
We compare the combination of Locret and HF-4bits quantization with the full attention method and the standalong HF-4bits quantization. We utilize the official implementation of Hugging Face, with Quanto as the backend of quantization. Other hyperparameters are kept same as described in Section 4.1. We conduct the experiment on L-Eval and report the average score, with Llama-3.1-8B-instruct backend. The results in Table 8 shows that the degradation caused by quantization is not significantly high, showing that Locret exhibits good robustness on data representation and it is friendly to quantization.
C.2 Combination with LoCoCo
Table 9: L-Eval scores of LoCoCo, Locret and the combination LoCoCo+Locret. (Detailed)
Llama-3.1-8B-instruct on L-Eval
Method CodeU NQ CUAD NarrativeQA QMSum SPACE Avg.$\uparrow$
FullAttn 10.0 66.84 38.91 23.11 18.76 16.86 29.08
LoCoCo 4.44 61.10 35.84 19.83 18.15 16.71 26.01
Locret 8.89 63.03 37.21 23.59 18.17 16.87 27.96
LoCoCo+Locret 7.78 66.33 38.01 24.85 18.31 16.92 28.70
We compare the combination of LoCoCo and Locret with the standalone methods. For LoCoCo, we train the convolution head with the size of convolved cache set to 2048. We extend the context length through chunked prefill training to 64K, which is longer than all tasks’ average input length. The convolution kernel is set to 21, and we train the newly-added convolution and layer norms for 200 steps, following the original setting. Since the original Llama-3.1-8B-instruct supports 128K context length, we do not modify its positional embedding. During Inference, we keep a cache budget size of 16384. In the standalone LoCoCo setting, there are 2048 cache units are convolved, while the others are the heavy-hitters selected by H2O. In the combined algorithm, we replace H2O to Locret. We select 14336 cache units with the highest CIS, and convolve the other evicted tokens into 2048 cache units. In all methods, we set the local length to 0, following the original setting.
Appendix D Training Robustness
Locret demonstrates high robustness to the training settings, suggesting that there is no need for careful tuning of training hyperparameters or meticulous selection of datasets. Here, we ablate the intermediate size of the retaining heads $d_{\mathbf{R}}$ and train the retaining head $\mathbf{R}$s on various long-context tuning datasets to demonstrate the stability of results across different training settings.
D.1 Intermediate Size of the retaining head
We align all the training settings as described in Section 4.1 and only change the intermediate size of retaining heads $d_{\mathbf{R}}\in\{256,512,1024,2048,4096\}$ with the backbone model Phi-3-mini-128K. The trained model is evaluated on L-Eval and we report the average L-Eval score corresponding to each intermediate size. Results are listed in Figure 2. The performance variations among all the settings are minimal compared to the changes in the intermediate size, surpassing all baselines in Table 2. This indicates that out method exhibits good performance stability regardless of the intermediate size of the retaining head $\mathbf{R}$s.
Table 10: L-Eval scores with different intermediate size of the retaining head $d_{\mathbf{R}}$. (Detailed)
Phi-3-mini-128K on L-Eval
$d_{\mathbf{R}}$ CodeU NQ CUAD NarrativeQA QMSum SPACE Avg.$\uparrow$
256 8.89 51.52 23.05 16.21 15.26 13.77 21.45
512 6.67 50.61 23.33 16.67 15.02 14.23 21.09
1024 8.89 51.49 22.23 16.42 14.86 14.06 21.33
2048 7.78 54.09 21.91 16.46 15.00 13.89 21.52
4096 10.00 52.33 23.52 16.15 14.81 14.02 21.81
We train different retaining head $\mathbf{R}$s with $d_{\mathbf{R}}\in\{256,512,1024,2048,4096\}$. We keep all the other hyperparameters same, and train on the same dataset. From Table 10, Locret shows stability to the intermediate size, in both overall performance and the performance of each single task. While increasing the intermediate size, we observe very slight overall performance enhancement. However, the performance variance is negligible compared to the increase of parameter size, thus we choose to maintain the intermediate size in a small scope to take balance of performance and efficiency.
D.2 Training Data Insensitivity
We also consider the sensitivity of the training data, which leads us to ablate the training dataset by training on LongAlign (Bai et al., 2024) and Anti-Haystack (Pan, 2024), comparing these results with those from LongAlpaca (Chen et al., 2024) in the original training setting. We also align other settings to the original setting and choose the backbone model to be Phi-3-mini-128K. We report the average L-Eval score for each training dataset. The results in Table 2 shows that Locret has high insensitivity towards different training data. The performance impact of different data recipes is minimal, indicating that our method can be trained on any long-context tuning dataset.
Table 11: L-Eval scores of Locret trained on various dataset. (Detailed)
Phi-3-mini-128K on L-Eval
Dataset CodeU NQ CUAD NarrativeQA QMSum SPACE Avg.$\uparrow$
LongAlpaca 8.89 51.49 22.23 16.42 14.86 14.06 21.33
LongAlign 10.00 55.13 21.34 16.40 15.01 14.09 22.00
Anti-Haystack 8.89 52.91 20.87 13.73 13.84 14.10 20.72
We conduct training on various datasets and benchmark the trained weights on L-Eval with Phi-3-mini-128K backbone, to show the stability towards training datasets. For each datasets, we set the training hyperparameters same and truncate the context to 10240 tokens. We train the first 3000 steps of LongAlpaca and LongAlign. Since Anti-Haystack is a relatively smaller dataset, we utilize the whole dataset, which consist of 2424 entries. The results in Table 11 shows that different training dataset recipe exhibits minor effect towards the overall performance. Locret can obtain competitive performance without delicately selecting the training data.
Appendix E Discontinuous Context and Stablizers
Evicting cache units results in context discontinuity, which causes unstable CIS prediction and inaccurate calculation of later tokens. Thus, we always retain the stabilizers, which are consist of the last $n_{s}$ cache units in each chunked prefill step. We ablate $n_{s}$ on R.Number of $\infty$-Bench in the proposed algorithm to demonstrate the necessity of incorporating stabilizers in the design. The results in Figure 3(a) show that lower stabilizer length $n_{s}$ causes severe performance degredation and the model fails completely when the stabilizers are absent. We report the maximum absolute error of the last hidden state of the input prompt across different layers in Figure 3(b). Large errors can be observed when the stabilizers are short or absent. We also report the mean absolute error of the predicted causal importance values with different stabilizer lengths, compared to the case without evicting any cache units, in Figure 3(c). We also observe high errors when the stabilizer length is limited. This explains the reason for failure when the stabilizers are short or absent: context discontinuity leads to instability in the prediction of CIS, resulting in errors during cache eviction and amplifying errors in the hidden states.
Appendix F Causal Importance Score Simulates a Cache Problem
In this section, we show that assigning each cache unit a CIS and calculate each cache units with top-$b$ cache units simulates a cache problem, i.e. the calculation process can be done in a cache. Thus, Locret mathmatically equals to top-$b$ sparse attention.
Definition F.1.
(Causal Calculation) Given a sequence of objects $c_{1},c_{2},\cdots,c_{n}$, if
$\forall 1\leq i\leq n$, $c_{i}=f(c_{1},c_{2},\cdots,c_{i-1})$
then $f$ is a causal calculation. $c_{1},c_{2},\cdots,c_{n}$ is the generated sequence respective to $f$.
For all causal calculations, we can easily split the function into two parts: a selection function and a another function. Formally,
[…]
$g:2^{\{c_{1},c_{2},\cdots,c_{n}\}}\rightarrow\{c_{1},c_{2},\cdots,c_{n}\},$
$Sel:2^{\{c_{1},c_{2},\cdots,c_{n}\}}\rightarrow 2^{\{c_{1},c_{2},\cdots,c_{n}\}};~{}X\mapsto Y\subseteq X,$
[…]
Definition F.2.
(Causal Importance Score) Given a causal calculation $f$ and $c_{1},c_{2},\cdots,c_{n}$ is the generated sequence of $f$. $s_{1},s_{2},\cdots,s_{n}\in\mathbb{R}$ is a sequence of numbers. If
$s_{i}=h(c_{i}),$
then $\{s_{i}\}$ is a CIS of sequence $\{c_{i}\}$. $h$ is a causal importance scoring function.
Definition F.3.
(Cache Problem) Given a causal calculation $f=g\circ Sel$, its generated sequence $\{c_{i}\}$ and a positive number $b\in\mathbb{Z}_{+}$, if $f$ satisfies the following two condion, then $(f,b,\{c_{i}\})$ is a cache problem with budget $b$.
  • $\forall 1\leq i\leq n,~{}|Sel(c_{1},\cdots,c_{n})|\leq b,$
  • $\forall 1\leq m_{1}<m_{2}\leq n,~{}Sel(c_{1},\cdots,c_{m_{2}})\backslash Sel(c_{1},\cdots,c_{m_{1}})\subseteq\{c_{m_{1}+1},\cdots,c_{m_{2}}\}.$
Theorem F.1.
(Calculating cache units with Top-$b$ CIS is a cache problem.) Given a causal calculation $f=g\circ Sel$, and its generated sequence $\{c_{i}\}$, a CIS $s_{i}=h(c_{i})$ and a positive number $b\in\mathbb{Z}_{+}$, if the selection function $Sel$ satisfies the following condition,
$Sel(c_{1},c_{2},\cdots,c_{i})=\{c_{p_{1}},c_{p_{2}}\cdots,c_{p_{b^{\prime}}}\},~{}s_{p_{1}},s_{p_{2}}\cdots,s_{p_{b^{\prime}}}\in\text{Top-}b(s_{1},s_{2},\cdots,s_{i})$
then $(f,b,\{c_{i}\})$ is a cache problem with budget $b$.
Proof. (1) For all $i$ of $1\leq i\leq n$, $|Sel(c_{1},\cdots,c_{i})|=|\{c_{p_{1}}\cdots,c_{p_{b^{\prime}}}\}|=|\{s_{p_{1}},\cdots,s_{p_{b^{\prime}}}\}|$. Since $s_{p_{1}},s_{p_{2}}\cdots,s_{p_{b^{\prime}}}\in\text{Top-}b(s_{1},s_{2},\cdots,s_{i})$, |$|\{s_{p_{1}},\cdots,s_{p_{b^{\prime}}}\}|\leq b$. Thus $|Sel(c_{1},\cdots,c_{i})|\leq b$.
(2) For all $1\leq m_{1},<m_{2},\leq n$,
$Sel(c_{1},\cdots,c_{m_{2}})\backslash Sel(c_{1},\cdots,c_{m_{1}})$$\subseteq\{c_{m_{1}+1},\cdots,c_{m_{2}}\}$
$\iff\{s_{p_{1}},\cdots,s_{p_{m_{2}}}\}\backslash\{s_{q_{1}},\cdots,s_{q_{m_{1}}}\}$$\subseteq\{s_{m_{1}+1},\cdots,s_{m_{2}}\}.$
Assume $\exists s\in\{s_{p_{1}},\cdots,s_{p_{m_{2}}}\}\backslash\{s_{q_{1}},\cdots,s_{q_{m_{1}}}\}$ but $s\notin\{s_{m_{1}+1},\cdots,s_{m_{2}}\}$. Since $s_{p_{1}},\cdots,s_{p_{m_{2}}}=\text{Top-}b(s_{1},\cdots,s_{m_{2}}),~{}s\in\{s_{1},\cdots,s_{m_{2}}\}$. Thus $s\in\{s_{1},\cdots,s_{m_{1}}\}$. $s$ is not in the Top-$b$ values of first $m_{1}$ scores, thus there exists $b$ values larger than $s$, denote as $s_{l_{1}},\cdots,s_{l_{b}}$. Then, $s_{p_{1}},\cdots,s_{p_{m_{2}}}=\text{Top-}b(s_{l_{1}},\cdots,s_{l_{b}},s_{m_{1}+1},\cdots,s_{m_{2}}).$ From this, we can obtain that $\min\{s_{p_{m_{2}}}\}\geq\min\{s_{l_{1}},\cdots,s_{l_{b}}\}>s$, $s\notin\{s_{p_{1}},\cdots,s_{p_{m_{2}}}\}$. Contradiction. Finally, there must be $s\in\{s_{m_{1}+1},\cdots,s_{m_{2}}\}$. From (1)(2), $f$ satisfies the two conditions of cache problem. Thus, calculating cache units with Top-$b$ CIS is a cache problem.
Appendix G Retained Patterns of Locret
[…]
Figure 6 shows that the pattern is mostly decided by the tasks, where both heads shows similar pattern in the same task. In R.Number, we are able to observe a strong signal between token 10000 and 15000, which is the position of the inserted number string, indicating that Locret can identify the potentially answer-related parts by giving high predicted values of CIS. In M.Find, we can observe the StreamingLLM (Xiao et al., 2024b) pattern, where the tokens at the beginning of the sequence are always important. This is also mentioned as the $\Lambda$-pattern in MInference. We can also discover the vertical lines in the middle of the sequence. This pattern is also approached by MInference (Jiang et al., 2024a) by the pattern “vertical-and-slash”. In E.MC, H2(Zhang et al., 2024e) and ScissorHands (Liu et al., 2024a) pattern can be observed, following the assumption that if a token is activated at some point, it will continue to be activated in the consequencing process. Noticing that the vertical lines always come in groups, which is the fundament of InfLLM (Xiao et al., 2024a) retrieving blocks to calculate. The comparison between two heads also shows that different heads exhibits different features. Head 22 of layer 11 shows stronger vertical lines at some point, where retained pattern of head 14 layer 11 is more even. Head 14 of layer 11 also gives stronger signal to the initial tokens, where this effect is less strong in head 22 layer 11. We also conduct experiments to investigate the patterns across layers. In Figure 6, we show that the pattern variance of the same head in different layers can be large. In shallow layers, e.g. layer 1 and 5, the retained cache units appears to be periodical and semantic independent. However, in middle layers, e.g. layer 13 and 17, the position of the inserted number string is strongly highlighted, indicating that semantic takes over to be the dominant factor. In the deepest layers, e.g. 21, 25 and 29, the highlighted vertical line at the position of the inserted string becomes more accurate.
The retained pattern at different layers shows various features, which might be a good handle to investigate how LLMs understand and process natural language queries.
Figure 5: Head patterns across multiple tasks.
Figure 6: Layer patterns of R.Number
Appendix H The Training Process of Locret
Here, we present changing trend of loss and accuracy during training in Figure 7.
Figure 7: Training loss and accuracy during the training process.

### v2

@@
-Locret: Enhancing Eviction in Long-Context LLM Inference with Trained Retaining Heads
+Locret: Enhancing Eviction in Long-Context LLM Inference with Trained Retaining Heads on Consumer-Grade Devices
 Abstract
-Large language models (LLMs) have shown remarkable advances in supporting long-context comprehension and processing tasks.
-However, scaling the generation inference of LLMs to such long contexts incurs significant additional computation load, and demands a substantial GPU memory footprint to maintain the key-value (KV) cache of transformer-based LLMs.
-Existing KV cache compression methods, such as quantization, face memory bottlenecks as context length increases, while static-sized caches, such as selective eviction, suffer from inefficient policies.
-These limitations restrict deployment on consumer-grade devices like a single Nvidia 4090 GPU.
-To overcome this, we propose Locret, an efficient framework for long-context LLM inference that introduces retaining heads to evaluate the causal importance of KV cache units, allowing for more accurate eviction within a fixed cache size.
-Locret is fine-tuned on top of the frozen backbone LLM using a minimal amount of data from standard long-context SFT datasets.
-During inference, we evict low-importance cache units along with a chunked prefill pattern, significantly reducing peak GPU memory usage.
-We conduct an extensive empirical study to evaluate Locret, where the experimental results show that Locret outperforms the recent popular and competitive approaches, including InfLLM, Quantization, SirLLM, and MInference, in terms of memory efficiency and the quality of generated contents — Locret achieves over a 20×20$\times$20 × and 8×8$\times$8 × KV cache compression ratio compared to the full KV cache for Phi-3-mini-128K and Llama-3.1-8B-instruct.
-Additionally, Locret can be combined with other efficient inference methods, such as quantization and token merging.
-To the best of our knowledge, Locret is the first framework capable of deploying Llama-3.1-8B or similar models on a single Nvidia 4090 GPU, enabling 128K long-context inference without compromising generation quality, and requiring little additional system optimizations.
-The experimental code and the tool kit of Locret is provided at https://github.com/huangyuxiang03/Locret
+Scaling the input context length of a large language model (LLM) incurs a significant increase in computation cost and memory footprint to maintain the attention key-value (KV) cache.
+Existing KV cache compression methods suffer from inefficient compression strategies and limited memory reduction effects, making it difficult for LLMs to conduct long-context inference on consumer-grade devices, especially when inferring long-context stream input.
+Such obstacles prevent consumer-grade devices from supporting more complex applications, creating challenges for the democratization of LLMs.
+To overcome this, we propose Locret, the first framework to create an eviction policy compatible with chunked prefill.
+By evaluating the causal importance of KV cache units by learnable retaining heads, Locret enables precise eviction of cache units, facilitating efficient long-context inference.
+In our extensive empirical studies, Locret outperforms the recent popular and competitive approaches in terms of memory efficiency and generation quality — Locret achieves up to 20×20$\times$20 × of KV cache compression ratio within less than $10\%$ performance loss.
+Furthermore, Locret achieves 128K+ long-context inference on a single NVIDIA 4090 GPU without compromising generation quality and only costs $<1$ GPU hour of additional training.
+We provide the code of Locret at https://github.com/huangyuxiang03/Locret.
+Long-context Inference, Memory Efficient Inference, Large Language Models
 1 Introduction
-Large language models (LLMs) have revolutionized AI development and deployment (Zhao et al., 2023; Minaee et al., 2024).
-Recent advancements in LLMs’ ability to handle long-context tasks have further unlocked the potential of generative AI.
-State-of-the-art LLMs now support significantly extended context lengths, with GPT-4o (OpenAI, 2024) and Llama-3.1 (Dubey et al., 2024) handling 128K tokens, Yi (Young et al., 2024) and Claude-3 (Anthropic, 2024) supporting 200K tokens, and Gemini-1.5 (Reid et al., 2024) reaching 10 million tokens.
-These advances enable LLMs to tackle complex tasks like multi-hop reasoning (Li et al., 2024; Schnitzler et al., 2024), solving needle-in-a-haystack problems (Guerreiro et al., 2023; Wang et al., 2024a), and powering advanced LLM agents (Qin et al., 2023; Wang et al., 2024b) and AI-driven operating systems (Mei et al., 2024).
-However, deploying generative inference under long-context settings on consumer-grade GPUs requires innovative algorithmic and system optimizations to handle this new paradigm efficiently.
-Compared to traditional short-context LLM inference, long-context LLM inference shifts the computing paradigm in two key ways: i) increased computational overhead for attention mechanisms: as context length grows, the computation required for obtaining attention scores increases quadratically, which results in a higher ratio of the computational budget in a transformer block; ii) higher memory footprint for key-value (KV) caching: longer contexts require larger KV caches, which dramatically increases the peak memory usage.
-These shifts demand innovative techniques to mitigate computational costs and manage memory usage effectively for long-context LLM inference.
-Although various efforts have been made to overcome the bottleneck in LLM inference, these approaches fail to enable long-context inference on consumer-grade GPUs.
-Models with compact architectures, e.g.
-MiniCPM-128K (Hu et al., 2024b) and Phi-3-mini-128K (Abdin et al., 2024), reduce the computational load and memory usage, but cannot alleviate the KV cache burden in long-context scenarios.
-Similarly, techniques like LLM model weight quantization (Frantar et al., 2023; Lin et al., 2024; Ma et al., 2024), activation quantization (Dettmers et al., 2022; Xiao et al., 2023; Zhang et al., 2024c), or sparsification (Liu et al., 2023; Zhang et al., 2022) also fall short in effectively reducing the memory usage to a level supported by consumer-grade GPUs.
-Recently, some specific optimizations have been proposed for long-context LLM inference.
-For example, sparse attention mechanisms (Jiang et al., 2024a; Ge et al., 2024; Lou et al., 2024) attempt to reduce runtime memory through conduct selected few calculation, and KV cache quantization (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024) reduces cache size by applying low-bit storage.
-These methods can only offer limited compression rate, of which the core issue is that the KV cache grows linearly with sequence length, and the above methods do not adequately address this problem.
-On the contrary, combining chunked prefill with token-dropping techniques (Xiao et al., 2024b; Yang et al., 2024) could offer a more effective solution, as it maintains a static-sized cache where the memory usage can be bounded.
-However, current token-dropping and cache eviction methods (Zhang et al., 2024e; Liu et al., 2024a; Yao et al., 2024), whose token importance manually designed according to the inner statistics, suffer from accuracy loss and performance degradation due to inaccuracies in estimating token importance — The weakening correlation between local and global importance as sequences grow exacerbates this issue.
-To address these limitations, we propose a lightweight training-based paradigm that provides more accurate token importance scoring to tackle the long-context LLM inference problem.
-We highlight our contributions below:
-Contribution 1: We propose a lightweight training-based paradigm for selective KV cache eviction for long-context LLM inference, with an offline training cost $<$1 GPU hours.
-We tackle the problem of KV cache eviction by a learning based approach.
-We introduce the retaining heads, with a small number of additional parameters, fine-tuned on top of the frozen backbone LLM using a minimal amount of data from standard long-context SFT datasets to estimate the causal importance of each cache unit.
-Such a training paradigm is able to provide accurate token importance scoring prediction and can be integrated with other efficient inference algorithms, e.g., quantization and token merging.
-Contribution 2: We provide an efficient inference system implementation for Locret.
-We integrate the retaining head mechanism into a chunked prefill inference framework, where we maintain a static-size cache set through evicting cache units with low predicted importance to limit the GPU memory usage.
-Locret is able to preserve the most important cache units with the trained retaining heads, enabling precise attention approximation without compromising the inference latency.
-Locret is also applicable to all transformer-based LLMs and various hardware, as it requires minimal modifications to the model’s inference process and only utilizes dense operators.
-Contribution 3: We conduct an extensive evaluation of Locret, which illustrates that Locret can not only obtain a comparable performance but also maintain inference efficiency.
-Locret achieves over a 20×20$\times$20 × and 8×8$\times$8 × KV cache compression ratio for Phi-3-mini-128K and Llama-3.1-8B-instruct, enabling full comprehension of long contexts on consumer-grade devices.
-To the best of our knowledge, Locret is the first framework capable of deploying Llama-3.1-8B or similar models on a single Nvidia 4090 GPU, enabling 128K long-context inference without compromising generation quality, and requiring little extra system optimizations.
+In recent years, large language models (LLMs) have revolutionized generative AI (Zhao et al., 2023; Minaee et al., 2024), and the advancements of LLMs in handling long-context tasks have further unlocked the potential of generative AI.
+As a result, the context lengths supported by state-of-the-art LLMs have been significantly extended, such as GPT-4o (OpenAI, 2024) handling 128K tokens, Claude-3 (Anthropic, 2024) supporting 200K tokens, and Gemini-1.5 (Reid et al., 2024) even reaching 10M tokens.
+These improvements enable LLMs to tackle complex applications with extremely long or streaming inputs, such as multi-hop reasoning (Li et al., 2024b; Schnitzler et al., 2024), LLM-driven agents (Qin et al., 2024b; Wang et al., 2024), and AI-powered operating systems (Mei et al., 2024).
+Some recent efforts (Hu et al., 2024b; Abdin et al., 2024) have successfully deployed LLMs on consumer-grade end-side devices instead of cloud servers and conducted inference with limited context.
+We envision that unleashing the potential of long-context inference on consumer-grade devices will revolutionize the development of personalized AI applications and the democratization of LLMs.
+However, conducting long-context LLM inference on consumer-grade devices remains a challenging problem that requires algorithmic innovations and systematic optimizations.
+Figure 1: For each prefix length of the context, this figure shows the consistency in evaluating the token importance of the prefix based on the full context and based on only the prefix without subsequent tokens.
+The consistency is defined as the intersection of the top 10% tokens of two evaluation methods divided by the number of top 10% tokens in the prefix.
+More details are in Appendix B.
+As context length scales, the challenge of long-context LLM inference arises from two major aspects: the increased computational cost of the attention mechanism and the higher memory footprint due to the key-value (KV) cache.
+This leads to the failure of traditional optimizations targeting model backbones to provide a sufficient solution.
+Specifically, backbone-targeted optimizations, such as compacting model architectures (Hu et al., 2024b; Abdin et al., 2024) and quantizing model weights (Frantar et al., 2023; Dettmers et al., 2022; Xiao et al., 2023; Lin et al., 2024), fail to improve the efficiency of attention patterns or the KV cache, as attention’s quadratic complexity with respect to sequence length remains unaddressed.
+To this end, recent efforts have focused on optimizing attention patterns and the KV cache to achieve efficient long-context LLM inference.
+Recent attention-targeting optimizations, including sparse attention (Jiang et al., 2024a; Ge et al., 2024; Lou et al., 2024) and KV cache quantization (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024), show promising results to accelerate attention computation and reduce memory footprint.
+However, they fail to fundamentally address the core challenge: the KV cache grows linearly with context length.
+Layer-wise chunked prefill combined with attention sparsity (e.g., SnapKV (Li et al., 2024a)) can alleviate this problem to a certain extent.
+This technique typically performs cache eviction after the precise attention computation for each layer and requires access to the entire sequence.
+It can theoretically support longer sequences by limiting the maximum memory usage to a single layer’s KV cache, but it cannot handle streaming input whose length grows continually.
+The combination of cache eviction methods (Xiao et al., 2024c; Yang et al., 2024) and chunked prefill offers a more promising approach by maintaining a static cache size and supporting streaming input.
+Yet, as shown in Figure 1, existing eviction techniques like H2O (Zhang et al., 2024d) and SnapKV show significant discrepancies between local importance estimation and global importance estimation, i.e.
+it is hard to estimate the importance of each token only based its previous tokens.
+Instead, these methods require a large number of subsequent tokens to make an accurate estimation.
+Other methods like SirLLM (Yao et al., 2024) show great local-global estimation consistency but suffer from performance degradation.
+To overcome these limitations, we propose a lightweight training-based paradigm, Locret, that provides a more accurate token importance estimation to select the victim during KV cache eviction, enabling efficient and scalable long-context LLM inference:
+Contribution 1: We propose Locret, a lightweight training-based paradigm for selective KV cache eviction in long-context LLM inference.
+It introduces learnable retaining heads to estimate the causal importance score (CIS) for token selection, with an offline training cost of $<$1 GPU hour.
+Additionally, we present Locret-Q, a query-aware variant of Locret, slightly modified to handle query-driven tasks (e.g., long document question answering).
+Contribution 2: We provide an efficient inference system implementation for Locret, integrating retaining heads into a chunked prefill framework.
+This integration limits GPU memory usage by evicting low-CIS KV cache units during the prefill process and accelerates the prefill time.
+Locret is compatible with all decoder-only LLMs and imposes minimal additional hardware requirements.
+Contribution 3: We extensively evaluate Locret, demonstrating its ability to achieve comparable performance with full KV cache while maintaining inference efficiency.
+Locret achieves over 20×20$\times$20 × and 8×8$\times$8 × KV cache compression ratios for Phi-3-mini-128K and Llama-3.1- 8B-instruct, respectively.
+Additionally, Locret-Q accelerates prefill by over 2×2$\times$2 × on query-driven tasks without significant performance degradation.
+This framework enables full comprehension of long contexts on consumer-grade devices without compromising the generation quality, and introduces minimal additional system optimizations.
 2 Related Work
-Efforts in long-context LLM inference can be categorized by algorithm and system optimizations:
-Algorithm optimizations.
-Optimizations aimed at reducing the size of the KV cache can generally be classified into three categories: quantization-based methods, token dropping methods, and sparsity-based methods.
-Quantization-based methods (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024; Zhang et al., 2024a), which store the KV cache in low-bit representations, require hardware support for these formats and may slow down inference due to the overhead of dequantization.
-Token dropping methods typically follow two main strategies: eviction or the use of an attention pool.
-Eviction-based approaches, such as H2(Zhang et al., 2024e), ScissorHands (Liu et al., 2024a), and SirLLM (Yao et al., 2024), rank tokens by certain statistical metrics to identify the most influential ones, discarding others to reduce memory usage.
-Attention pool-based methods (Nawrot et al., 2024; Rajput et al., 2024), such as StreamingLLM (Xiao et al., 2024b) and LoCoCo (Cai et al., 2024), compress multiple adjacent KV cache units into a single unit using a specially designed transformation.
-Sparsity-based methods (Ge et al., 2024; Jiang et al., 2024a; Yang et al., 2024; Lou et al., 2024; Lv et al., 2024) focus on leveraging the sparsity patterns of attention heads to reduce both computation and I/O.
-The combination of these approaches can be further enhanced by identifying specific attention patterns for each head and layer (Ge et al., 2024).
-For surveys of these methods, please refer to (Yuan et al., 2024; Kang et al., 2024; Shi et al., 2024).
-System optimizations.
-The challenge of long-context inference can also be alleviated from a system-level perspective.
-Offloading-based methods (Sheng et al., 2023; Xiao et al., 2024a; Wu et al., 2024; Lee et al., 2024) store the KV cache in CPU memory, retrieving only the most relevant chunks to the GPU before computing a new chunk.
-This approach reduces peak GPU memory usage, though at the cost of slower inference.
-Hardware-aware algorithms, such as flash attention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) and page attention (Kwon et al., 2023), exploit GPU architecture (Ghorpade et al., 2012) to enable more efficient runtime memory management.
-In addition, reimplementing inference infra-structure in a more efficient programming language (llama.cpp, ; llama2.c, ; rustformers, ), or adopting disaggregated inference (Jiang et al., 2024b; Zhong et al., 2024; Qin et al., 2024; Hu et al., 2024a), can greatly enhance inference efficiency.
-Algorithmic optimizations can be seamlessly integrated into such systems (Agrawal et al., 2023; Lee et al., 2024).
-For instance, KTransformers (KVCache.AI, 2024) adopts the chunked offloading technique from InfLLM (Xiao et al., 2024a).
-However, system optimizations primarily focus on extending context length by leveraging hardware resources, rather than directly reducing the size of the KV cache.
-3 Locret
-3.1 Preliminaries
-Transformer architecture.
-We define the model inference of transformer-based LLMs as follows.
-Given a token sequence $t_{1},t_{2},\cdots,t_{n}$, we denote the output hidden state of layer $i$ as $\textbf{H}^{(i)}$ and $\textbf{H}^{(0)}$ is the embeddings.
-Each transformer layer consists of an attention layer and an MLP layer.
-We assume the model follows a grouped-query attention (GQA) architecture (Ainslie et al., 2023), with $h$ query heads and a group size of $g$.
-For multi-head attention (MHA), $g$ is set to 1.
-The attention score of layer $i$’s head $j$ is calculated by $\mathbf{A}_{j}^{(i)}=\text{softmax}\left(\mathbf{Q}_{j}^{(i)}{\mathbf{K}_{\lceil{j/g}\rceil}^{(i)T}}/\sqrt{d_{m}}\right)\cdot\mathbf{V}_{\lceil{j/g}\rceil}^{(i)}$, where $d_{m}$ represents the hidden size for each head and $\left[\mathbf{Q}_{j}^{(i)},\mathbf{K}_{j}^{(i)},\mathbf{V}_{j}^{(i)}\right]=\mathbf{H}^{(i-1)}\cdot\left[{\mathbf{W}_{j}^{Q(i)}},{\mathbf{W}_{j}^{K(i)}},{\mathbf{W}_{j}^{V(i)}}\right]$.
-Next, we compute $\mathbf{A}^{(i)}=\left[\mathbf{A}_{1}^{(i)},\cdots,\mathbf{A}_{h}^{(i)}\right]\cdot\mathbf{W}^{O(i)}$, finally followed by $\mathbf{H}^{(i)}=\text{MLP}(\mathbf{A}^{(i)})$.
-KV cache and chunked prefill.
-During the prefill stage, all prompt tokens are processed in a single forward pass, where $\mathbf{Q}^{(i)}$, $\mathbf{K}^{(i)}$, and $\mathbf{V}^{(i)}$ each have a sequence length of $n$.
-In the decoding stage, only a single token is processed across layers, utilizing the KV cache units to reduce computation.
-Chunked prefill is a method for reducing peak memory consumption by processing tokens in chunks over multiple passes, with the assistance of the KV cache.
-Taking both KV cache and chunked prefill into account, the attention calculation can be modified as Equation 1, where $B$ represents the number of tokens processed in a single model pass.
-For decoding, $B=1$, while for chunked prefill, $B$ corresponds to the chunk size.
-We denote the attention output for tokens $n+1,n+2,\cdots,n+B$ as $\mathbf{A}[n+1\colon n+B]$, and the attention output for the $k$-th token as $\mathbf{A}[k]$.
-$\mathbf{A}{[n+1\colon n+B]}_{j}^{(i)}=\text{softmax}\left(\frac{\mathbf{Q}{[n+1\colon n+B]}_{j}^{(i)}{\mathbf{K}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)T}}}{\sqrt{d_{m}}}\right)\cdot\mathbf{V}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)}.$ (1)
-Cache eviction.
-Cache eviction in long-context inference is defined as follows.
-Here, we slightly abuse the notation of heads and layers, and treat the key-value vector pair of a single token within one head as the smallest cache unit.
-We denote the cache unit for the $k$-th token as $c_{k}=(\mathbf{K}[k],\mathbf{V}[k])$.
-Assume a memory budget $b$, representing the maximum number of cache units that can be stored in GPU memory at any given time.
-The abstract form of attention can then be written as $c_{k}=f(c_{1},c_{2},\cdots,c_{k-1})$.
-With limited cache capacity, this calculation can only be approximated by $\tilde{c}_{k}=f(\tilde{c}_{p_{1}},\tilde{c}_{p_{2}},\cdots,\tilde{c}_{p_{b^{\prime}}})$, where $b^{\prime}\leq b$, and $p_{1},p_{2}\cdots,p_{b^{\prime}}\in\{1,2,\cdots,k-1\}$.
-Intuitively, the number of prior cache units involved cannot exceed the memory budget.
-When $b^{\prime}=b$, indicating the cache is full, one cache unit must be evicted.
-We select the unit to be evicted using a policy $p_{v}=\text{Policy}(\tilde{c}_{p_{1}},\cdots,\tilde{c}_{p_{b}};\tilde{c}_{k})$.
-In such stated problem, the key challenge of cache eviction is to develop an effective policy function that minimizes the approximation error $\|\tilde{c}_{k}-c_{k}\|$.
+This paper focuses on optimizing long-context LLM inference.
+Existing efforts can be categorized into algorithm and system optimization.
+For more details about LLMs, please refer to the surveys (Zhao et al., 2023; Lu et al., 2024).
+Algorithm Optimizations aim to reduce the size of the KV cache and can generally be classified into three categories: quantization-based methods, sparsity-based methods, and token dropping methods.
+Quantization-based methods (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024; Zhang et al., 2024a) use low-bit values to represent the KV cache, reducing cache memory overhead and improving cache computing efficiency.
+These quantization-based methods suffer from hardware-oriented operator customization and additional inverse quantization overhead.
+Sparsity-based methods (Ge et al., 2024; Jiang et al., 2024a; Yang et al., 2024; Lou et al., 2024; Lv et al., 2024) leverage the sparsity patterns of attention heads to reduce both computational and I/O costs.
+Combining different patterns can yield further optimization by identifying specific patterns for each head (Ge et al., 2024; Jiang et al., 2024a; Xiao et al., 2024b).
+For more details on sparsity-based methods, please refer to the surveys (Yuan et al., 2024; Kang et al., 2024; Shi et al., 2024).
+Although quantization-based methods and sparsity-based methods have achieved promising results, they cannot address the issue that the KV cache memory overhead increases linearly with the context length.
+Eviction-based methods, such as H2O (Zhang et al., 2024d), ScissorHands (Liu et al., 2024a), and SirLLM (Yao et al., 2024), rank KV cache units by certain statistical metrics to identify the most influential units, discarding others to reduce memory usage and speed up attention computation.
+Pooling-based methods (Nawrot et al., 2024; Rajput et al., 2024), especially StreamingLLM (Xiao et al., 2024c) and LoCoCo (Cai et al., 2024a), compress multiple adjacent KV cache units into a single unit using pre-designed transformations.
+More important units will merge into compressed units with higher weights.
+Eviction-based and pooling-based methods drop or merge tokens to maintain a static cache size, but struggle with accurate victim selection and optimal pooling function design.
+System Optimizations alleviate the challenge of long-context inference from a system-level perspective, by fully considering hardware features.
+Offloading-based methods (Sheng et al., 2023; Xiao et al., 2024a; Wu et al., 2024; Sun et al., 2024) use CPU memory to store the KV cache and retrieve only the most relevant chunks to GPU memory before inferring a new chunk.
+These methods reduce maximum GPU memory usage at the cost of introducing CPU-GPU communication overhead.
+Hardware-aware methods, such as Flash-Atttention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) and Page-Attention (Kwon et al., 2023), enable more efficient runtime memory management by considering GPU architectures (Ghorpade et al., 2012).
+In addition, building inference infrastructures with a more efficient programming language (llama.cpp, ; llama2.c, ; rustformers, ), or adopting disaggregated inference (Jiang et al., 2024b; Zhong et al., 2024; Qin et al., 2024a; Hu et al., 2024a), can also greatly improve long-context inference efficiency.
+Since system optimizations primarily enhance efficiency by leveraging hardware resources rather than directly optimizing attention patterns or the KV cache, relying solely on them cannot adequately address the challenges of long-context LLM inference.
+Several efforts have integrated algorithm optimizations into system optimizations (Agrawal et al., 2023; Lee et al., 2024), such as KTransformers (KVCache.AI, 2024) leveraging offloading based on InfLLM (Xiao et al., 2024a), achieving promising results.
 Figure 2: The framework of Locret.
@@
 “t” represents the time step in chunked prefill, “$b$” represents the budget size, and “$n_{s}$” represents the length of the stabilizers.
-3.2 Locret  Framework
-Locret is a training-based KV cache compression framework that works in conjunction with chunked prefill.
-As illustrated in Figure 1, Locret operates in two stages: training and inference.
-In the training stage, we modify the original LLM by appending a retaining head $\mathbf{R}$ to each attention module.
-We then train the retaining heads $\mathbf{R}$ while keeping the LLM backbone frozen.
-During chunked prefill inference, the retaining heads $\mathbf{R}$ are used to calculate the importance of each cache unit in the chunk.
-We retain the cache units with higher scores, along with stabilizers (i.e., the last tokens), in the cache pool located in GPU memory.
-Through this process, the retaining heads $\mathbf{R}$ learn and predict the patterns discovered by existing methods, as detailed in Appendix G.
-From a mathematical perspective, cache eviction is performed by assigning each cache unit an importance score that reflects its influence on the calculation of subsequent tokens.
-We refer to this estimation as the causal importance score (CIS) since it is computed in a causal manner.
-The CIS of cache unit $k$ is calculated as $s_{k}=S(c_{1},c_{2},\cdots,c_{k})$.
-By applying top-$b$ sparse attention based on the CIS, we can ensure that the trace (i.e.
-retaining and eviction) of each cache unit can be fit within a cache with a given memory budget.
-Further details can be found in Appendix F.
-However, since not all tokens can be stored in the cache simultaneously, calculating the actual CIS on-chip is impractical.
-Instead, we use a heuristic approximation for CIS, defined as follows: $\tilde{s}_{k}=S(\tilde{c}_{p_{1}},\tilde{c}_{p_{2}},\cdots,\tilde{c}_{p_{b^{\prime}}})$, where $b$ is the cache budget, all $\tilde{c}_{p_{i}}$ are approximated cached units, and $b^{\prime}\leq b$.
-We hypothesize that if the scoring function for causal importance is sufficiently accurate, it will consistently select the most critical cache units, resulting in a negligible difference between the heuristic and the actual score.
-Thus, we use the terms heuristic CIS and actual CIS interchangeably.
-3.3 Training the Retaining Heads
-In this section, we introduce Locret’s model architecture modifications and the corresponding training recipe.
-We add additional parameters to compute the CIS $s_{k}$ (or $\tilde{s}_{k}$ for on-chip inference) with respect to all previous cache units.
-Specifically, we inject a retaining head, consisting of a small MLP, into each layer.
-The retaining head predicts the CIS for each head of the corresponding layer based on the concatenation of $[\mathbf{Q},\mathbf{K},\mathbf{V}]$.
-Formally, with a slight abuse of notation, let the retaining head for layer $i$ be denoted as $\mathbf{R}$.
-The CIS at head $j$ of layer $i$ is then calculated as: $\mathbf{\tilde{S}}=\mathbf{R}([\mathbf{Q},\mathbf{K},\mathbf{V}])=\sigma([\mathbf{Q},\mathbf{K},\mathbf{V}]\mathbf{W}_{1})\mathbf{W}_{2}$.
-In this equation, $\mathbf{W}_{1}\in\mathbb{R}^{(d_{m}+2d_{kv})\times d_{\mathbf{R}}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{\mathbf{R}}\times\frac{h}{g}}$ are the tunable parameters of $\mathbf{R}$, $\sigma$ is the activation function and $\mathbf{\tilde{S}}[k]_{j}$ is the predicted CIS of the $k$-th token at head $j$ of layer $i$.
-This architecture implies that the importance estimation for a single head is not performed in isolation but rather considers all heads together.
-Note that for GQA models, there are only $h/g$ output values corresponding to the number of heads in the KV cache.
-We train the retaining head $\mathbf{R}$s on a small Question-Answer (QA) supervised fine-tuning (SFT) dataset, where each entry consists of a single prompt and one answer.
-We define the CIS $s_{k}$ for the $k$-th token as the maximum attention score, before softmax, from all the answer tokens toward the $k$-th token.
-Formally, for the $k$-th token at head $j$ of layer $i$, we approximate the predicted value $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ to the ground truth $\mathbf{S}[k]_{j}^{(i)}:=\max_{p}\left(\mathbf{Q}_{j}^{(i)}\mathbf{K}_{j}^{(i)T}\right)_{p,k}$, where $n_{q}(d)\leq p\leq n_{q}(d)+n_{a}(d)$, and $n_{q}(d)$ and $n_{a}(d)$ represent the lengths of the prompt and answer in data $d$, respectively.
-For an MHA model with $L$ layers and $h$ heads, the training objective is described in Equation 2.
-For GQA models, we take the maximum attention score before softmax across different query heads within the same group as the ground truth for the corresponding KV head.
-$\operatorname*{arg\,min}_{\mathbf{W}_{1}^{(i)},\mathbf{W}_{2}^{(i)},i=1,2\cdots,L}\mathbb{E}_{d\in\mathcal{D}}\left[\sum_{i=1}^{L}\sum_{j=1}^{h}\sum_{k=1}^{n_{q}(d)}\mathcal{L}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)\right]$ (2)
-The training loss consists of a regression loss and a smoothing loss.
-We apply the Smooth-$\mathcal{L}_{1}$ norm between the predicted values and the ground truth.
-Since important segments in natural language typically consist of adjacent tokens, we also apply the $\mathcal{L}_{2}$ norm between each pair of adjacent predicted values to enforce smoothness.
-The complete training loss for Locret is given by Equation 3.
-$\mathcal{L}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)=\text{Smooth-}\mathcal{L}_{1}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)+\alpha\mathcal{L}_{2}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{\tilde{S}}[k+1]_{j}^{(i)}\right)$ (3)
-Figure 2: L-Eval scores with different intermediate size of retaining head $d_{\mathbf{R}}$.
-Table 1: L-Eval scores of Locret trained on various datasets.
-Dataset LongAlpaca LongAlign Anti-Haystack
-L-Eval 21.33 22.00 20.72
-From our observations, the training of Locret exhibits strong robustness.
-The performance variations shown in Figure 2 and Table 2 are minimal, despite changes in $d_{\mathbf{R}}$ and the dataset.
-Details can be found in Appendix D.
-Training statistics, including loss dynamics, are recorded in Appendix H.
-3.4 Inference Implementation of Locret
-Input: Model $\mathbf{M}$, Prompt tokens $x$, Local length $n_{loc}$, Stablizer length $n_{s}$, Budget $b$, Chunk size $B$
-Output: Generated tokens $x_{gen}$
-// Leave the last $n_{loc}$ out to make sure they are not evicted.
-chunk_positions $\leftarrow$ split_chunk(00, $x$.length() $-n_{loc}$, $B$)
-K_cache, V_cache, score_cache $\leftarrow$ [], [], []
-for chunk $\in$ chunk_positions do
-      begin_pos, end_pos $\leftarrow$ chunk.begin_pos, chunk.end_pos
-[…]
-      K_cache $\leftarrow$ Concat(K_cache, K_chunk)
-      V_cache $\leftarrow$ Concat(V_cache, V_chunk)
-      score_cache $\leftarrow$ Concat(score_cache, score_chunk)
-      if chunk is not the last chunk then
-           // Keep the last $n_{s}$ caches to maintain higher context continuity.
-           score_cache[score_cache.length()-$n_{s}$:score_cache.length()] $\leftarrow$+$\infty$ + ∞
-           end if
-          indices $\leftarrow$ top-$b$(score_cache).indices
-           K_cache, V_cache, score_cache = K_cache[indices], V_cache[indices], score_cache[indices]
-           end for
-          K_cache, V_cache, score_cache $\leftarrow\mathbf{M}$($x$[$x$.length()$-n_{loc}$:$x$.length()], K_cache, V_cache)
-           $x_{gen}\leftarrow\mathbf{M}$.generate(K_cache, V_cache)
-           return $x_{gen}$
-Algorithm 1 Locret Inference
-During the inference stage, we use the chunked prefill pattern and perform cache eviction based on the predicted CIS.
-Since the predicted value $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ depends solely on $\mathbf{Q}[k]_{j}^{(i)}$, $\mathbf{K}[k]_{j}^{(i)}$, and $\mathbf{V}[k]_{j}^{(i)}$, and because attention in decoder-only models is causal, $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ remains consistent once calculated.
-Thus, we store the KV cache units along with their corresponding causal importance values.
-When the cache is full, we evict the units with lower causal importance values, as they are deemed less useful for future computations.
-Cache eviction introduces context discontinuity, meaning some cache units at certain positions may be absent.
-This can degrade generation quality and increase the error between the predicted and accurate CIS, as LLMs are typically not trained on such contexts.
-To mitigate this, we retain the last tokens of the current chunk at each step of the chunked prefill process, ensuring a local and continuous context to minimize errors.
-To demonstrate the effectiveness of this design, we perform an ablation study on the length of stabilizers $n_{s}$, shown in Figure 3.
-Smaller $n_{s}$ results in severe performance degradation, and the model fails entirely when stabilizers are absent, as context discontinuity leads to instability in CIS prediction, causing errors in cache eviction and amplifying errors in hidden states.
-More details are discussed in Appendix E.
+For simplicity, our notation here does not reflect the concept of layers.
+3 Methodology of Locret
+3.1 Preliminaries
+Transformer Architecture.
+Given a token sequence $\{t_{1},\cdots,t_{n}\}$ as the input prompt of the transformer-based LLM, we denote the output hidden states of the $i$-th layer as $\textbf{H}^{(i)}$, and denote the input embeddings of the first layer as $\textbf{H}^{(0)}$.
+For each transformer layer, it consists of an attention block and a feedforward neural network (FFN) block.
+Attention blocks often follow the grouped-query attention (GQA) architecture (Ainslie et al., 2023), with $h$ query heads and $h/g$ KV heads, where $g$ is the group size, i.e., $g$ heads share the same KV heads.
+The multi-head attention (MHA) architecture adopted in the original transformer can be regarded as a special GQA ($g=1$).
+In the $i$-th layer, the attention score of the $j$-th query head is formalized as $\mathbf{A}_{j}^{(i)}=\texttt{softmax}\left(\mathbf{Q}_{j}^{(i)}{\mathbf{K}_{\lceil{j/g}\rceil}^{(i)\top}}/\sqrt{d_{m}}\right)\cdot\mathbf{V}_{\lceil{j/g}\rceil}^{(i)}$ , where $d_{m}$ represents the hidden size for each head and $\left[\mathbf{Q}_{j}^{(i)},\mathbf{K}_{\lceil{j/g}\rceil}^{(i)},\mathbf{V}_{\lceil{j/g}\rceil}^{(i)}\right]=\mathbf{H}^{(i-1)}\cdot\left[{\mathbf{W}_{j}^{(i)\texttt{Q}}},{\mathbf{W}_{\lceil{j/g}\rceil}^{(i)\texttt{K}}},{\mathbf{W}_{\lceil{j/g}\rceil}^{(i)\texttt{V}}}\right]$ After obtaining the attention score, the output of the $i$-th attention block is $\mathbf{A}^{(i)}=\left[\mathbf{A}_{1}^{(i)},\cdots,\mathbf{A}_{h}^{(i)}\right]\cdot\mathbf{W}^{(i)\texttt{O}}$ , and the output hidden states of the $i$-th layer is $\mathbf{H}^{(i)}=\texttt{FFN}(\mathbf{A}^{(i)})$
+KV Cache and Chunked Prefill.
+Given the input prompt sequence $\{t_{1},\cdots,t_{n}\}$, during the prefill stage, all prompt tokens are processed in a single forward pass.
+After the prefill, $\mathbf{K}^{(i)}=\left[\mathbf{K}^{(i)}_{1},\cdots,\mathbf{K}^{(i)}_{h/g}\right]$ and $\mathbf{Q}^{(i)}=\left[\mathbf{Q}^{(i)}_{1},\cdots,\mathbf{Q}^{(i)}_{h/g}\right]$ are stored as the KV cache, whose sequence length is $n$.
+During the decoding stage, each time a token is decoded, a forward pass is conducted only for this token and decode the next token.
+In this process, the KV cache is used to avoid redundant attention computation.
+Chunked prefill is a method for reducing peak memory usage by spliting tokens into mulitple chunks and prefilling tokens chunk by chunk.
+Taking both the KV cache and chunked prefill into account, the attention block can be modified as follows,
+$(\displaystyle\mathbf{A}{[n+1\colon n+B]}_{j}^{(i)}=\texttt{softmax}\Biggl{(}$1)
+$\frac{\mathbf{Q}{[n+1\colon n+B]}_{j}^{(i)}\mathbf{K}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)\top}}{{\sqrt{d_{m}}}}\Biggr{)}$ $\mathbf{V}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)},$
+where $\mathbf{A}[n+1\colon n+B]$ denotes the attention output for the tokens $\{n+1,\cdots,n+B\}$, and $B$ is the number of tokens processed in a single forward pass.
+For decoding, $B=1$, while for chunked prefill, $B$ corresponds to the chunk size.
+For the $k$-th token in the context, its attention output is $\mathbf{A}[k]$, its key and value vectors are $\mathbf{K}[k]$ and $\mathbf{V}[k]$.
+Cache Eviction.
+In the cache eviction process, we treat the KV vector pair of a single token within one attention head as the smallest cache unit.
+We denote the cache unit of the $k$-th token as $c_{k}=(\mathbf{K}[k],\mathbf{V}[k])$.
+Assuming a memory budget $b$, representing the maximum number of cache units that can be stored, the abstract form of the attention block can be written as $c_{k}=f(c_{1},c_{2},\cdots,c_{k-1})$.
+With limited cache capacity, this process can only be approximated by $\tilde{c}_{k}=f(\tilde{c}_{p_{1}},\tilde{c}_{p_{2}},\cdots,\tilde{c}_{p_{b^{\prime}}})$, where $b^{\prime}\leq b$, and $p_{1},p_{2}\cdots,p_{b^{\prime}}\in\{1,2,\cdots,k-1\}$.
+When the cache is full, one unit must be evicted.
+We select the victim using some policy $p_{v}=\text{Policy}(\tilde{c}_{p_{1}},\cdots,\tilde{c}_{p_{b}};\tilde{c}_{k})$, and the key challenge is to develop an effective policy that minimizes the error $\|\tilde{c}_{k}-c_{k}\|$.
 Figure 3: R.Number with different stabilizer lengths $n_{s}$.
@@
 We conduct this experiment on entries 101-120 of R.Number using the Phi-3-mini-128K backbone.
-We maintain a cache pool with a capacity of $b$ cache units, discarding units that exceed this limit.
-For each chunk, the model processes the chunked input tokens alongside the current cache pool.
-The newly generated KV pairs and their predicted scores are then concatenated with the existing cache.
-Once the cache pool is full, only the $b$ cache units with the highest CIS values are retained.
-At each chunked prefill step, except for the final step, we retain the stabilizers, i.e.
-the last $n_{s}$ cache units.
-Additionally, we do not compress the last $n_{loc}$ tokens of the prompt, as they are critical for maintaining high generation quality due to their strong correlation with the query.
-Finally, the answer is generated according to the compressed KV cache.
-Algorithm 1 provides the pseudocode for Locret, where we formally describe the Locret inference process.
-The GPU memory usage during Locret inference can be effectively bounded.
-GPU memory for KV cache storage is limited to $\mathcal{O}(b+n_{loc})$, and the runtime memory usage of the attention mechanism is bounded by $\mathcal{O}(B\times(b+B+n_{loc}))$.
-For comparison, while processing an input with $n$ tokens, full attention prefill requires $\mathcal{O}(n)$ for KV cache storage and $\mathcal{O}(n^{2})$ for runtime memory, whereas chunked prefill requires $\mathcal{O}(n)$ for KV cache storage and $\mathcal{O}(nB)$ for runtime memory consumption.
+3.2 Framework of Locret
+Locret is a training-based KV cache eviction framework that works in conjunction with chunked prefill.
+As illustrated in Figure 2, Locret operates in two stages: training and inference.
+In the training stage, we modify the original LLM by appending a retaining head $\mathbf{R}$ to each attention module.
+We then train the retaining heads $\mathbf{R}$ while keeping the LLM backbone frozen.
+During the chunked prefill inference stage, the retaining heads $\mathbf{R}$ can obtain the importance of each cache unit.
+We retain the cache units with higher scores, along with stabilizers (i.e., the last tokens), in the cache pool located in GPU memory.
+Through this process, the retaining heads $\mathbf{R}$ learn and predict the patterns discovered by existing methods, e.g.
+attention sink in Xiao et al.
+(2024c) and vertical lines in MInference (Jiang et al., 2024a), as detailed in Appendix O.
+The eviction policy assigns each cache unit an importance score reflecting its influence on comprehending subsequent context.
+This estimation is causal, termed the causal importance score (CIS).
+The CIS of the $k$-th unit depends only on the preceding units and the $k$-th unit itself.
+Due to memory constraints, calculating the exact CIS on-chip is impractical.
+Intuitively, discarding certain KV cache during the attention calculation introduces errors, manifesting as biases in the CIS.
+However, as long as the importance estimation is sufficiently accurate, the core information can be preserved, and the loss of information in the attention process can be minimized.
+Therefore, the error introduced in the CIS can be neglected.
+Please refer to Appendix N for more details.
+3.3 Training Retaining Heads
+In this section, we introduce Locret’s model architecture modifications and the corresponding training recipe.
+We add additional parameters to compute the CIS $\mathbf{S}[k]$ for the $k$-th cache unit.
+Specifically, we inject a retaining head, consisting of a small MLP, into each layer.
+From our observation, such small MLPs do not slow down model inference, with details elaborated in Appendix M.
+The retaining head predicts the CIS for each head of the corresponding layer based on the concatenation of $[\mathbf{Q},\mathbf{K},\mathbf{V}]$.
+Formally, with a slight abuse of notation, let the retaining head for layer $i$ be denoted as $\mathbf{R}$.
+The CIS at head $j$ is then calculated as: $\mathbf{\tilde{S}}=\mathbf{R}([\mathbf{Q},\mathbf{K},\mathbf{V}])=\sigma([\mathbf{Q},\mathbf{K},\mathbf{V}]\mathbf{W}_{1})\mathbf{W}_{2}$ In this equation, $\mathbf{W}_{1}\in\mathbb{R}^{(d_{m}+2d_{kv})\times d_{\mathbf{R}}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{\mathbf{R}}\times\frac{h}{g}}$ are the tunable parameters of $\mathbf{R}$, and $\sigma$ is the activation function, $\mathbf{\tilde{S}}[k]=\left[\mathbf{\tilde{S}}[k]_{1},\cdots,\mathbf{\tilde{S}}[k]_{h/g}\right]\in\mathbb{R}^{\frac{h}{g}}$ , where $\mathbf{\tilde{S}}[k]_{j}$ is the predicted CIS of the $k$-th token.
+This architecture implies that the importance estimation for a single head is not performed in isolation but considers all heads.
+We train the retaining head $\mathbf{R}$s on a small Question-Answer (QA) supervised fine-tuning (SFT) dataset, where each entry consists of a single prompt and one answer.
+We define the CIS $\mathbf{S}[k]_{j}$ for the $k$-th token at head $j$ as the maximum attention score, before softmax, from all the answer tokens toward the $k$-th token.
+Formally, given a training instance $d$, for the $k$-th token at head $j$ of layer $i$, we approximate the predicted value $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ to the ground truth $\mathbf{S}[k]_{j}^{(i)}:=\max_{p}\left(\mathbf{Q}_{j}^{(i)}\mathbf{K}_{j}^{(i)T}\right)_{p,k}$ , where $n_{q}(d)<p\leq n_{q}(d)+n_{a}(d)$, and $n_{q}(d)$ and $n_{a}(d)$ represent the lengths of the prompt and answer in data $d$, respectively.
+For an MHA model with $L$ layers and $h$ heads, the training objective is described in Equation 2.
+For GQA models, we take the maximum attention score before softmax across different query heads within the same group as the ground truth for the corresponding KV head.
+$\underset{\mathbf{W}_{1}^{(i)},\mathbf{W}_{2}^{(i)},i=1,2\cdots,L}{\text{argmin}}\mathbb{E}_{d\in\mathcal{D}}\left[\sum_{i=1}^{L}\sum_{j=1}^{h}\sum_{k=1}^{n_{q}(d)}\mathcal{L}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)\right]$ (2)
+The training loss consists of a regression loss and a smoothing loss.
+We apply the Smooth-$\mathcal{L}_{1}$ norm between the predicted values and the ground truth.
+Since important segments in natural language typically consist of adjacent tokens, we also apply the $\mathcal{L}_{2}$ norm between each pair of adjacent predicted values to enforce smoothness.
+The complete training loss for Locret is given by Equation 3.
+[…]
+$+\alpha\mathcal{L}_{2}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{\tilde{S}}[k+1]_{j}^{(i)}\right)$
+From our observations, the training of Locret exhibits strong robustness.
+Despite changes in $d_{\mathbf{R}}$ and the dataset, the performance variations shown in Figure 6 and Table 6 are minimal.
+Details can be found in Appendix I.
+Training statistics and loss dynamics are recorded in Appendix P.
+3.4 Inference Implementation of Locret
+During the inference stage, we use the chunked prefill pattern and perform cache eviction based on the predicted CIS.
+Since the predicted CIS does not rely on subsequent tokens, it remains consistent once calculated.
+Thus, we store the KV cache units along with their corresponding causal importance values.
+When the cache is full, we evict the units with lower causal importance values, as they are deemed less useful for future computations.
+Such eviction is performed during chunked prefill.
+When processing a new chunk, we first compute its KV cache, concatenate it with the previously retained cache, and evict redundant units to adhere to the budget size.
+Note that we cache the pre-RoPE KV cache and reassign continuous position embeddings from the beginning to enhance context continuity.
+Cache eviction introduces context discontinuity, meaning some cache units at certain positions may be absent, which can degrade generation quality.
+To mitigate this, we retain the last $n_{s}$ tokens of the current chunk, named as the stabilizers, at each step of chunked prefill, ensuring a local and continuous context to minimize errors.
+As shown in Figure 3, smaller $n_{s}$ results in severe performance degradation, and the model fails entirely when stabilizers are absent, as context discontinuity leads to instability in CIS prediction, causing errors in cache eviction and amplifying errors in hidden states.
+More details are discussed in Appendix L.
+We provide a pseudocode of Locret inference in Algorithm 1.
 4 Experiments
-In this section, we present the experiments conducted to evaluate the proposed framework, Locret, aiming to address the following questions:
-(Q1) Can Locret obtain better end-to-end task performance compared to popular and competitive long-conetext inference approaches within similar or less peak memory?
-(Q2) Can Locret improve inference speed compared to other approaches?
-(Q3) Can Locret be combined with other efficient inference methods?
-4.1 Experimental Setup
-Models and training dataset.
-We evaluate Locret on two long-context LLMs: Phi-3-mini-128K (Abdin et al., 2024) and Llama-3.1-8B-instruct (Dubey et al., 2024).
-Both models can process up to 128K context tokens and follow MHA and GQA architectures, respectively.
-Phi-3-mini-128K has 3.8B parameters, while Llama-3.1-8B-instruct contains 8B parameters, making both suitable for deployment on consumer-grade devices.
-We inject retaining heads $\mathbf{R}$ into each layer, setting the intermediate size $d_{\mathbf{R}}$ to 1024 for both models.
-The retaining heads are trained on the LongAlpaca dataset (Chen et al., 2024) for 3000 steps.
-Training Locret is lightweight, with the tunable parameters comprising 8% and 2.5% of the total for the two models, respectively.
-The complete training process takes 0.47 and 0.80 GPU hours on a single A800 GPU for each corresponding model.
-The cache budget $b$ is set to 6000 for Phi-3-mini-128K and 16384 for Llama-3.1-8B-instruct.
-More details on hyperparameters for both training and inference, as well as the system environment, can be found in Appendix A.
-[…]
-Baselines.
-As discussed in Section 2, existing algorithms for memory-efficient long-context inference can be categorized into offloading-based, quantization-based, token-dropping, and sparsity-based methods.
-For each category, we select one representative method as the baseline.
-We compare Locret against full attention inference (denoted as FullAttn), InfLLM (Xiao et al., 2024a), KV cache quantization (Turganbay, 2024), SirLLM (Yao et al., 2024), and MInference (Jiang et al., 2024a).
-For quantization, we use Hugging Face Quanto (Hugging-Face, ) implementation, referring to the 2-bit quantization method as HF-2bits.
-We omit HF-4bits and benchmark this combination in Section 4.4.
-We do not include attention pool-based token-dropping methods in this benchmark, as they are orthogonal to our approach; further discussion is provided in Section 4.4.
-Detailed introductions to the selected baselines can be found in Appendix A.
-We also discuss the comparison between the trained Locret and the randomly initialized retaining heads $\mathbf{R}$ in Appendix B.
-4.2 End-to-end Benchmark
-[…]
-[…]
-In the L-Eval tests, all methods show some degree of performance degradation.
-Nevertheless, Locret achieves the best overall result, obtaining the highest score on most tasks.
-L-Eval is a shorter but more complex dataset, where SirLLM performs particularly well.
-Quantization fails on most tasks, resulting in the lowest overall score.
-Both InfLLM and MInference suffer significant performance drops compared to FullAttn inference.
-Locret consistently surpasses all competitors.
-[…]
-Figure 4: Memory Statistics vs.
-Task Performance.
-The red lines correspond to the theoretical size of the model weights, while the blue lines represent the total theoretical size of the model weights and the full KV cache without any compression.
-The purple lines indicate the accuracies of FullAttn.
-“Total Memory” represents the total memory usage of both GPU and CPU memory.
+We conduct experiments to evaluate whether Locret can address the following questions:  (Q1) Can Locret obtain better end-to-end task performance compared to popular and competitive long-context inference methods using similar or less peak memory?
+(Q2) Is Locret able to achieve a faster inference speed on consumer-grade devices?
+(Q3) How can Locret process query-driven tasks?
 […]
@@
 Locret 8.89 63.03 37.21 23.59 18.17 16.87 27.96
-4.3 Speed Test on Real Consumer-Grade Devices
+Figure 4: Memory Statistics vs.
+Task Performance.
+The red lines correspond to the theoretical size of the model weights, while the blue lines represent the total size of the model weights and the full KV cache without any compression.
+The purple lines indicate the accuracies of FullAttn.
+“Total Memory” represents the total memory usage of both GPU and CPU.
+4.1 Experimental Setup
+Model and Dataset Settings.
+We conduct experiments on two long-context LLMs: Phi-3-mini-128K (Abdin et al., 2024) and Llama-3.1-8B-instruct (Dubey et al., 2024).
+Both models can process up to 128K context tokens and follow MHA and GQA architectures, respectively.
+The parameter sizes of these two models are also suitable for deployment on consumer-grade devices.
+We inject retaining heads $\mathbf{R}$ into each layer of these two models, and the intermediate size $d_{\mathbf{R}}$ is 1,024.
+The retaining heads are trained on the LongAlpaca dataset (Chen et al., 2024) for 3,000 steps , with a 5e-4 learning rate, 10,240 sequence length, and $\alpha$ set to 2.5e-3.
+Training Locret is lightweight, with the tunable parameters comprising 8% and 2.5% of the total for the two models, respectively.
+The complete training process takes 0.47 and 0.80 GPU hours on an A800 GPU for each corresponding model.
+More important hyperparameters are listed in Table 4.
+More details on hyperparameters and system environments can be found in Appendix A.
 […]
-R.PassKey is a task where the model retrieves a 5-digit number from a large amount of irrelevant text, a task we believe to be relatively simple for humans.
-Thus, we consider the task to have failed if the accuracy falls below 95%.
-As shown in Table 3, aside from the settings that fail on this task, Locret achieves the highest inference speed among all methods that can correctly process R.PassKey.
-Due to its MHA architecture, Phi-3-mini-128K has a larger KV cache, which leads to failures for both HF-2bits and MInference.
-Storing the full KV cache on a single 4090 GPU is infeasible, as it requires 48GB of memory.
-Although the quantized KV cache is reduced to 6GB, the converting processes between representations requires significant GPU memory for intermediate states, resulting in the failure of HF-2bits.
-While InfLLM can run in memory-limited scenarios, its offloading process slows down inference, with I/O becoming the bottleneck in attention calculation.
-SirLLM fails due to its inaccurate eviction policy, which cannot correctly identify the 5-digit number.
-In the GQA model (Llama-3.1-8B-instruct), which has a smaller KV cache, the quantized cache can fit within the GPU memory.
-However, the quantization and dequantization processes become the bottleneck, leading to significantly slower speeds.
-The performance of InfLLM, SirLLM, and MInference is similar to that seen with Phi-3-mini-128K.
-Although MInference benefits from faster encoding speeds, it fails on this task because it cannot process the entire input sequence at once.
-Locret strikes a balance between inference speed and performance, making it a far more suitable solution for long-context scenarios on consumer-grade devices.
+Baselines.
+As discussed in Section 2, existing algorithms for memory-efficient long-context inference can be categorized into offloading-based, sparsity-based, quantization-based, and token-dropping methods.
+For each category, we select one representative method as the baseline.
+We compare Locret against full attention inference (denoted as FullAttn), InfLLM (Xiao et al., 2024a), MInference (Jiang et al., 2024a), KV cache quantization (Turganbay, 2024), and SirLLM (Yao et al., 2024).
+For quantization, we use HuggingFace Quanto (Hugging-Face, ) implementation, referring to the 2-bit quantization method as HF-2bits.
+We omit HF-4bits and benchmark the combination with Locret in Section H.
+We do not include attention pooling-based token-dropping methods in this benchmark, as they are orthogonal to our approach, and further discussion about this is provided in Section H.
+Detailed introductions to the selected baselines can be found in Appendix A.
+We also discuss the comparison between the trained Locret and the randomly initialized retaining heads $\mathbf{R}$ in Appendix E.
+4.2 End-to-end Benchmark
+[…]
+[…]
+(2) On L-Eval, all methods show performance degradation.
+Nevertheless, Locret achieves the best overall performance, obtaining the highest scores on most tasks.
+L-Eval is a shorter but more complex dataset, where SirLLM performs particularly well.
+Quantization fails on most tasks.
+Both InfLLM and MInference suffer significant performance drops compared to FullAttn inference.
+Locret consistently surpasses all competitors.
+We report memory consumption in Figure 4, showing:
+[…]
+(2) For Phi-3-mini-128K, which has a larger KV cache, InfLLM and MInference exhibit higher memory consumption since they store the full KV cache.
+Although other methods have similar memory footprints, Locret uses the least memory and achieves the best overall performance.
+(3) For Llama-3.1-8B-instruct, whose full KV cache is smaller, the memory bottleneck shifts to the runtime computational memory for attention and other calculations.
+All methods exhibit similar memory footprints, with Locret delivering the best overall performance.
+In summary, our experiments demonstrate that Locret is both effective and efficient, outperforming all baselines on multiple datasets and models while using less GPU memory.
 Table 2: Executing R.PassKey on an NVIDIA 4090.
 “tok/s” represents the inference speed, “C.Len” stands for the context length after truncation, and “Acc.” represents task accuracy.
-“Success” shows whether a setting has $\geq$95accuracy.
 The highest score among 128K context is marked in bold.
@@
 Acc.$\uparrow$ OOM 99.83 OOM 1.69 OOM 100.00 0.00 13.56
-Success
 Llama-3.1- 8B-instruct tok/$\uparrow$ - 2287.66 1365.51 1589.75 - 3209.10 3680.06 5135.74
@@
 Acc.$\uparrow$ OOM 100.00 35.59 1.69 OOM 100.00 26.78 20.34
-Success
-4.4 Orthogonality to Other Methods
-We evaluate the combination of Locret with quantization and token merging to address Q3.
-Table 4: Quantization with FullAttn and Locret. “M” represents Method and “$-\Delta$” represents the gap of average L-Eval score.
-Setting M M-4bits $-\Delta$
-M=FullAttn 29.08 28.52 0.56
-M=Locret 27.96 27.11 0.85
-Table 5: The average L-Eval scores of LoCoCo, Locret, and the combination of LoCoCo and Locret.
-Method LoCoCo Locret Combination
-L-Eval 26.01 27.96 28.70
-KV cache quantization.
-According to Zhang et al.
-(2024d), eviction-based methods like H2O struggle with compatibility when combined with KV cache quantization.
-Quantization introduces significant disturbance in the estimation of heavy-hitters, leading to severe performance degradation.
-However, Locret is not affected by such issues and can be combined with quantization while maintaining most of its performance.
-Here, we compare the performance degradation caused by quantization on Locret with that of the full attention method using the same metrics.
-We use Quanto as the quantization backend and report the average L-Eval score with Llama-3.1-8B-instruct as the model backbone.
-Table 4.4 shows that the performance drop caused by quantization on Locret is only slightly higher than that observed with the full attention method, indicating that Locret is a quantization-friendly approach.
-More details of the experiment are provided in Appendix C.1.
-Token merging.
-As described in Section 2, token dropping can also be implemented through an attention pool.
-Attention pool-based methods (Xiao et al., 2024b; Cai et al., 2024; Mu et al., 2024; Munkhdalai et al., 2024) merge adjacent tokens or cache units into an attention pool, maintaining a static cache size.
-These methods are orthogonal to Locret , as the evicted tokens can be merged into a small cache pool and retained in GPU memory.
-We conduct the following experiment to demonstrate that Locret can serve as an effective plug-in scoring function within such frameworks, enhancing performance without increasing memory budget.
-We select LoCoCo (Cai et al., 2024) as a representative of the latest attention pool-based methods.
-LoCoCo maintains a cache set consisting of two parts: the heavy hitters and the convolved non-heavy hitters.
-During each chunked prefill step, LoCoCo first identifies a set of heavy hitters according to H2(Zhang et al., 2024e), then applies 1-D convolution to the non-heavy hitters to compress them into a static size.
-By replacing H2O’s heavy-hitter scoring function with Locret, we retain the cache units with high CIS and convolve the others.
-We compare this combination with standalone LoCoCo and Locret on L-Eval using the Llama-3.1-8B-instruct backbone and report the average score across all selected tasks.
-As shown in Table 4.4, Locret achieves a higher score than LoCoCo, and the combined algorithm outperforms both standalone methods.
-This suggests that Locret provides a more accurate scoring function compared to H2O, and the two methods complement each other, demonstrating their orthogonality.
-Further details of the experiment are provided in Appendix C.2.
+4.3 Processing Speed on Real Consumer-Grade Devices
+[…]
+R.PassKey is a task where the model retrieves a 5-digit number from a large amount of irrelevant text, a task we believe to be relatively simple for humans.
+Thus, we consider the task to have failed if the accuracy falls below 95%.
+As shown in Table 2, aside from the settings that fail on this task, Locret achieves the highest inference speed among all methods that can correctly process R.PassKey:
+(1) Due to its MHA architecture, Phi-3-mini-128K has a larger KV cache, which leads to failures for both HF-2bits and MInference.
+Storing the full KV cache on a single 4090 GPU is infeasible, as it requires 48GB of memory.
+Although the quantized KV cache is reduced to 6GB, converting representations requires significant GPU memory for its intermediate states, resulting in the failure of HF-2bits.
+While InfLLM can run in memory-limited scenarios, its offloading process slows down inference.
+SirLLM fails due to its inaccurate eviction, which cannot correctly identify the 5-digit number.
+(2) In the GQA model (Llama-3.1-8B-instruct), which has a smaller KV cache, the quantized cache can fit within the GPU memory.
+However, the quantization and dequantization processes become the bottleneck, leading to significantly slower speed performance.
+The performance of InfLLM, SirLLM, and MInference is similar to that seen with Phi-3-mini-128K.
+Although MInference benefits from faster encoding speed, it fails this task because it cannot process the entire input sequence at once.
+Locret strikes a balance between inference speed and performance, making it a far more suitable solution for long-context scenarios on consumer-grade devices.
+4.4 Locret-Q: Supporting Query-Driven Tasks
+We mainly address question Q3 in this section.
+Query-driven tasks are characterized by highly sparse yet query-correlated critical regions within the context.
+A representative example is the Multikey-NIAH task from RULER (Hsieh et al., 2024), where the context comprises JSON key-value pairs given by UUIDs, and the objective is to retrieve specific values using given keys.
+Such contexts are inherently challenging to compress effectively without query information.
+As introduced by Sun et al.
+(2024), existing eviction-based techniques exhibit significant performance degradation when applied to query-driven tasks.
+Due to this, we evaluate Locret against selected eviction-based baselines using the RULER benchmark (Hsieh et al., 2024).
+We introduce Locret-Q, a query-aware variant of Locret.
+When training the retaining heads, we prepend the last $l_{q}$ query tokens to the sequence and gather CIS labels.
+At inference, the query is inserted at the sequence start, ensuring its visibility across all eviction actions.
+This adaptation enables Locret-Q to perform query-aware eviction.
+Table 3 shows Locret-Q against SnapKV (Li et al., 2024a), H2O (Zhang et al., 2024d), and SirLLM (Yao et al., 2024) on RULER with a 128K context length.
+For reference, we also include results for FullAttn and MInference.
+Metrics include prefill speed, decode speed, and task performances are reported.
+As shown in Table 3, all efficient inference methods exhibit performance degradation to some extent.
+However, Locret-Q outperforms other eviction-based methods and even surpasses MInference, demonstrating its effectiveness on query-driven tasks.
+SnapKV shows performance degradation, while H2O and SirLLM fail completely on RULER.
+For speedup, methods combining eviction and chunked prefill (Locret-Q, Locret, and SirLLM) significantly reduce prefill time, achieving >2×>2$\times$> 2 × speedup over FullAttn.
+SnapKV cannot accelerate prefill due to no computation reduction.
+H2O suffers from extremely slow prefill as it relies on full-sequence attention scores, incompatible with efficient implementations like Flash-Attention.
+Decoding speeds are similar across eviction-based methods, as they maintain comparable KV cache sizes after eviction, and all of them are faster than FullAttn.
+Notably, Locret fails on RULER, showing a gap compared to Locret-Q, highlighting the necessity of query-awareness for query-centric tasks.
+A simple modification to Locret unlocks its potential for such tasks.
+Table 3: Performance, prefill speed, and decode speed on RULER.
+The best and second-highest scores among eviction-based methods in each column are highlighted in bold and underlined.
+“” indicates only testing the first 20 entries per subtask due to poor performance.
+FullAttn is implemented using Flash-Attention.
+Method RULER-128K (%) Prefill (tok/s) Decode (tok/s)
+FullAttn 82.20 4319.95 12.32
+MInference 72.97 7205.06 2.61
+Locret-Q 75.54 9587.84 40.11
+SnapKV 48.76 4203.34 36.49
+H2O 15.04 464.73 44.70
+SirLLM 13.23 9717.41 40.86
+Locret 34.33 9587.09 37.38
+4.5 Additional Experiments
+Additional experiments that further investigate Locret are included in the appendices due to space limitations.
+We compare Locret with randomized eviction in Appendix E, evaluate Locret on LongBench in Appendix F, provide a hyperparameter analysis in Appendix G, and explore the combination of Locret with other methods in Appendix H.
+Please refer to the appendices for further details.
 5 Conclusion
-We propose Locret, a lightweight training-based method that enables memory-efficient inference of long contexts on consumer-grade devices.
-Locret introduces a few additional parameters (retaining heads) to predict the CIS of each cache unit during chunked prefill and performs cache eviction based on the predicted CIS.
-We conduct extensive experiments across different models and multiple datasets to compare Locret with major efficient inference techniques, including offloading, quantization, eviction, and sparse attention.
-Results show that Locret outperforms all baselines, using less GPU memory and without requiring offloading to CPU memory.
-Furthermore, Locret can be combined with other efficient inference methods, such as KV cache quantization and attention pool-based token dropping.
-Locret maintains high inference throughput on consumer-grade devices and demonstrates robustness to variations in the intermediate size of retaining heads and training data.
-The complete framework of Locret, including both training and inference, highlights its suitability for low-resource computing scenarios.
-Locret can be applied to various application scenarios, such as end-side multi-modal model inference and context compression during disaggregated inference, which we plan to explore in future work.
-We also aim to test Locret on additional model architectures, including multi-latent models and models of varying sizes.
-Additionally, we are highly interested in integrating Locret into end-side deployment infra-structures to enhance inference efficiency, particularly for devices like smartphones and personal computers.
-Acknowledgment
-[…]
+We propose Locret, a lightweight training-based method that enables memory-efficient long-context LLM inference on consumer-grade devices.
+Locret introduces retaining heads to predict the CIS of each cache unit during chunked prefill and performs accurate cache eviction.
+We conduct extensive experiments across different models and multiple datasets to compare Locret with major efficient inference techniques, and results show that Locret outperforms all baselines, using less GPU memory and without requiring offloading to CPU memory.
+Locret-Q, a query-aware variant of Locret, can further process query-centric tasks without significant performance degradation.
+Future work will involve testing Locret on other model architectures, such as encoder-decoder and multi-latent models.
+More evaluations of Locret on other popular devices like NVIDIA Jetson are also planned.
+We also plan to explore integrating existing KV cache budget allocation methods with Locret to further enhance inference efficiency.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Hyperparameters, Environment and Baselines
@@
 A.2 Inference
-Table 6: Hyperparameters in Locret’s inference stage.
+Table 4: Hyperparameters in Locret’s inference stage. “$b$” is cache budget, “$B$” refers to chunk size of chunked prefill, “$n_{s}$” refers to stabilizers length and “$n_{loc}$” is local length.
 Model $b$ $B$ $n_{s}$ $n_{loc}$
@@
 For all the experiments except the 4090 experiments in Section 4.3, we use a workstation with 8$\times$×NVIDIA A800/H800 GPUs and 104 Intel(R) Xeon(R) Platinum 8470 CPUs.
+We only use 1 GPU from the cluster for training, as the GPU requirements are less than 80GB for all training procedures.
 The device has 1.0 TB CPU memory.
@@
 SirLLM is an eviction-based token dropping algorithm, where tokens with low token-entropy is evicted once the cache is fullfilled.
+We use the official implementation of SirLLM, which includes some CPU operations including importance sorting.
 MInference is a typical method of reducing peak GPU memory consumption through rule-based sparse attention, but it does not reduce the size of KV cache.
@@
 We only use the short factor of RoPE for InfLLM, and no further model modification is conducted for HF-2bits and SirLLM.
-Appendix B The Effect of Training
+Appendix B The Global and Local Discrepancy of Scoring Functions
+Cache importance scoring functions can generally be categorized into two types: causal and non-causal.
+Non-causal functions, e.g.
+H2O and SnapKV, require information from subsequent cache units to determine the importance score of a cache unit, making them dependent on prefilling the entire sequence.
+On the other hand, causal functions, e.g.
+SirLLM and Locret, predict cache importance without relying on subsequent information.
+Non-causal scoring functions are incompatible with chunked prefill because they cannot calculate scores without access to the full sequence.
+If such functions are integrated with chunked prefill, they often face a significant discrepancy between the local importance score (without considering subsequent information) and the global importance score (with full context).
+To investigate this discrepancy, we measure the consistency of the top 10% most important cache positions identified in prefixes of various lengths compared to the full context.
+For reference, the full context is truncated to 6K tokens.
+The results shown in Figure 1 highlights that scoring functions requiring future information, such as H2O and SnapKV, suffer from significant discrepancies when subsequent cache units are not considered.
+SirLLM, while also causal, shows notable inaccuracies, leading to performance degradation as demonstrated in Table 1 and Table 2.
+[…]
+[…]
+[…]
+Method R.Number E.Sum E.MC C.Debug Avg.$\uparrow$
+FullAttn 97.12 17.92 55.46 23.10 48.40
+H2O 3.39 15.35 45.41 20.57 21.18
+SnapKV 2.54 15.44 41.92 21.43 20.33
+Locret 97.46 16.82 46.29 29.71 47.57
+Appendix C Pseudocode of Locret
+We provide the pseudocode of Locretin this section, and we describe the inferece process of Locret in Algorithm 1.
+Algorithm 1 Locret Inference
+  Input:Model $\mathbf{M}$, Prompt tokens $x$, Local length $n_{loc}$, Stablizer length $n_{s}$, Budget $b$, Chunk size $B$
+  Output:Generated tokens $x_{gen}$
+  // Leave the last $n_{loc}$ out to make sure they are not evicted.
+  chunk_positions $\leftarrow$ split_chunk(00, $x$.length() $-n_{loc}$, $B$)
+  K_cache, V_cache, score_cache $\leftarrow$ [], [], []
+  for chunk $\in$ chunk_positions do
+    begin_pos, end_pos $\leftarrow$ chunk.begin_pos, chunk.end_pos
+    K_chunk, V_chunk, score_chunk $\leftarrow$
+[…]
+    K_cache $\leftarrow$ Concat(K_cache, K_chunk)
+    V_cache $\leftarrow$ Concat(V_cache, V_chunk)
+    score_cache $\leftarrow$ Concat(score_cache, score_chunk)
+    if chunk is not the last chunk then
+       // Keep the last $n_{s}$ caches to maintain higher context continuity.
+       score_cache[score_cache.length()-
+       $n_{s}$:score_cache.length()] $\leftarrow$+$\infty$ + ∞
+    end if
+    indices $\leftarrow$ top-$b$(score_cache).indices
+    K_cache, V_cache, score_cache = K_cache[indices],
+    V_cache[indices], score_cache[indices]
+  end for
+  K_cache, V_cache, score_cache $\leftarrow$
+  $\mathbf{M}$($x$[$x$.length()$-n_{loc}$:$x$.length()], K_cache, V_cache)
+  $x_{gen}\leftarrow\mathbf{M}$.generate(K_cache, V_cache)
+  return $x_{gen}$
+Appendix D Evaluating Locret-Q on RULER
+Table 6: Performance, prefill speed, and decode speed on RULER (Detailed).
+The best and second-highest scores among eviction-based methods in each column are highlighted in bold and underlined, respectively.
+“” indicates testing limited to the first 20 entries per subtask due to poor performance.
+FullAttn is implemented using Flash-Attention.
+Method SG1 SG2 SG3 MK1 MK2 MK3 MV MQ VT CWE FWE QA1 QA2 Avg.
+FullAttn 99.40 99.80 99.60 98.20 87.60 67.00 94.65 98.00 60.98 71.40 72.20 78.20 41.6 82.20
+MInference 100.00 98.60 99.00 95.40 58.20 23.80 84.35 95.70 66.40 45.94 74.67 67.80 38.80 72.97
+Locret-Q 100.00 99.80 99.60 75.00 98.80 98.60 66.95 85.50 52.64 30.90 80.27 53.20 40.80 75.54
+SnapKV 100.00 82.80 11.60 70.40 6.20 1.40 68.50 78.40 49.04 31.90 50.07 51.20 32.40 48.76
+H2O 20.00 0.00 0.00 0.00 0.00 0.00 3.75 2.50 1.00 0.00 53.33 100.00 15.00 15.04
+SirLLM 0.00 5.00 5.00 0.00 5.00 5.00 6.25 8.75 12.00 0.00 80.00 30.00 15.00 13.23
+Locret 100.00 45.00 35.00 10.00 5.00 0.00 20.00 17.50 69.00 46.50 73.33 20.00 5.00 34.33
+To evaluate Locret-Q’s performance on query-centric tasks, we compare it with selected eviction-based baselines: SnapKV, H2O, SirLLM, and vanilla Locret.
+We also include FullAttn (implemented with Flash-Attention) and MInference for reference.
+The RULER benchmark consists of 500 synthetic queries per task, each with a context length of 128K tokens.
+All methods are tested on Llama-3.1-8B-instruct.
+For Locret-Q and Locret, we set the budget size $b$ to 6000, chunk size $B$ to 4096, stabilizers length $n_{s}$ to 2500, and local length $n_{loc}$ to 100.
+For SnapKV, the voting window size is set to 100, with the last 100 tokens retained.
+For H2O, due to its reliance on full-sequence attention scores, we use a layer-wise chunked prefill pattern with a chunk size of 1024.
+Larger chunk size would result in an out-of-memory error.
+For SirLLM, we configure the start size to 4, recent size to 1000, and budget size to 6000.
+All evaluations are conducted on a single NVIDIA A800-80GB GPU.
+For prefill and decode speed testing, all methods except H2O are implemented with Flash-Attention; H2O uses PyTorch’s vanilla attention due to its incompatibility with efficient attention implementations.
+Appendix E Trained Retaining Heads vs. Random Eviction
 […]
@@
 […]
-Appendix C Orthogonality to Other Methods
-C.1 Combination with Quantization
+Appendix F Evaluation on LongBench
+We conduct additional experiments to evaluate Locret on LongBench (Bai et al., 2024b), comparing it with baselines such as Full Attention, MInference, InfLLM, and SirLLM.
+For this evaluation, we used Phi-3-mini-128K with a retained head trained on LongAlign.
+To ensure a fair comparison, we excluded all Chinese subtasks from LongBench and focused solely on the English subtasks, as Phi-3-mini-128K was not specifically trained on Chinese corpora.
+The results are presented below.
+For Locret , we follow the hyperparameters presented in Table 4.
+Table 8: LongBench scores of Locret compared with baselines.
+Method gov_ report triviaqa narrative qa qmsum musique 2wikimqa multifield qa_en repobench -p qasper hotpotqa multi_ news trec passage_ retrieval_en passage _count samsum lcc Avg.$\uparrow$
+FullAttn 33.35 86.38 18.21 19.51 19.82 33.37 49.82 58.02 41.07 43.06 26.57 67.00 93.50 2.97 23.15 51.86 41.73
+MInference 32.94 86.87 19.46 19.57 18.85 33.30 49.14 58.98 40.31 43.56 26.35 68.00 89.00 2.10 25.58 53.68 41.73
+SirLLM 32.92 85.61 21.08 21.59 24.32 34.97 48.52 59.15 40.17 47.00 26.44 65.50 63.00 3.00 23.11 51.83 40.51
+InfLLM 25.96 84.87 20.83 19.61 13.63 27.43 41.29 55.73 30.51 38.05 25.36 64.50 10.00 7.50 0.28 61.59 32.95
+Locret 33.46 82.39 24.56 23.35 25.12 35.93 52.77 57.16 40.17 48.70 26.41 62.00 83.00 3.00 26.37 52.61 42.31
+We also report the maximum memory usage, including the GPU memory, the CPU memory, and the total maximum memory, alongside the average score on LongBench.
+For FullAttn, we exclude the maximum memory usage, aligning with Figure 4.
+Table 9: Comparison of methods on LongBench and memory usage.
+Method LongBench Max GPU Memory Max CPU Memory Total Max Memory
+FullAttn 41.73 - - -
+MInference 41.73 27.63 0.17 27.80
+SirLLM 40.51 18.29 0.05 18.34
+InfLLM 32.95 20.03 8.95 28.98
+Locret 42.31 17.71 0.15 17.86
+From the experiments above, Locret demonstrates the best overall performance and excels in the majority of subtasks.
+It outperforms all the baselines without any noticeable performance degradation while consuming less memory.
+Although MInference also avoids performance drops, it requires more GPU memory compared to Locret.
+SirLLM achieves comparable memory usage but shows some performance decline compared to FullAttn and Locret.
+InfLLM exhibits the most significant performance drop, and its offloading mechanism results in the highest CPU memory usage, making it the method with the largest total memory consumption.
+These results highlight Locret as an outstanding approach for evaluation on LongBench.
+Appendix G Hyperparameter Analysis
+We examine three key hyperparameters: budget, stabilizer length, and chunk size.
+Budget To evaluate the robustness of Locret under different budget constraints, we compare the proposed method with SnapKV (Li et al., 2024a) using chunked prefill on LongBench (Bai et al., 2024b).
+As shown in Figure 5(a), when the budget size increases, Locret demonstrates a faster performance improvement compared to SnapKV.
+Stabilizers Length As discussed in Figure 3, stabilizers play a crucial role in context retrieval tasks.
+However, in NLU tasks, the stability of $n_{s}$ remains relatively high.
+We evaluate QMSum with different stabilizer lengths $n_{s}$, with the budget set at 6000.
+As illustrated in Figure 5(b), performance remains consistent when $n_{s}$ is small.
+The observed performance degradation at larger $n_{s}$ values is due to the reduced space available for other cache units.
+Chunk Size Executing long-context inference on hardware with varying GPU memory limitations choices of chunk size.
+When the chunk size changes, Locret shows stable performance.
+We test on the NQ dataset from L-Eval using multiple chunk sizes ranging from 256 to 4096.
+The results, shown in Figure 5(c), highlight the stability of $n_{s}$.
+Figure 5: Scores of Locret under (a) various budgets; (b) various $n_{s}$; (c) various chunk size.
+Appendix H Orthogonality to Other Methods
+Table 10: Quantization with FullAttn and Locret. “M” represents Method and “$-\Delta$” represents the gap of average L-Eval score.
+Setting M M-4bits $-\Delta$
+M=FullAttn 29.08 28.52 0.56
+M=Locret 27.96 27.11 0.85
+Table 11: The average L-Eval scores of LoCoCo, Locret, and the combination of LoCoCo and Locret.
+Method LoCoCo Locret Combination
+L-Eval 26.01 27.96 28.70
+KV cache quantization.
+According to Zhang et al.
+(2024c), eviction-based methods like H2O struggle with compatibility when combined with KV cache quantization.
+Quantization introduces significant disturbance in the estimation of heavy-hitters, leading to severe performance degradation.
+However, Locret is not affected by such issues and can be combined with quantization while maintaining most of its performance.
+Here, we compare the performance degradation caused by quantization on Locret with that of the full attention method using the same metrics.
+We use Quanto as the quantization backend and report the average L-Eval score with Llama-3.1-8B-instruct as the model backbone.
+Table H shows that the performance drop caused by quantization on Locret is only slightly higher than that observed with the full attention method, indicating that Locret is a quantization-friendly approach.
+More details of the experiment are provided in Appendix H.1.
+Token merging.
+As described in Section 2, token dropping can also be implemented through an attention pool.
+Attention pool-based methods (Xiao et al., 2024c; Cai et al., 2024a; Mu et al., 2024; Munkhdalai et al., 2024) merge adjacent tokens or cache units into an attention pool, maintaining a static cache size.
+These methods are orthogonal to Locret , as the evicted tokens can be merged into a small cache pool and retained in GPU memory.
+We conduct the following experiment to demonstrate that Locret can serve as an effective plug-in scoring function within such frameworks, enhancing performance without increasing memory budget.
+We select LoCoCo (Cai et al., 2024a) as a representative of the latest attention pool-based methods.
+LoCoCo maintains a cache set consisting of two parts: the heavy hitters and the convolved non-heavy hitters.
+During each chunked prefill step, LoCoCo first identifies a set of heavy hitters according to H2O (Zhang et al., 2024d), then applies 1-D convolution to the non-heavy hitters to compress them into a static size.
+By replacing H2O’s heavy-hitter scoring function with Locret, we retain the cache units with high CIS and convolve the others.
+We compare this combination with standalone LoCoCo and Locret on L-Eval using the Llama-3.1-8B-instruct backbone and report the average score across all selected tasks.
+As shown in Table H, Locret achieves a higher score than LoCoCo, and the combined algorithm outperforms both standalone methods.
+This suggests that Locret provides a more accurate scoring function compared to H2O, and the two methods complement each other, demonstrating their orthogonality.
+Further details of the experiment are provided in Appendix H.2.
+[…]
+[…]
+[…]
+Method R.Number E.Sum E.MC C.Debug Avg.$\uparrow$
+Locret 97.46 16.82 46.29 29.71 47.57
+Locret+PyramidKV 99.66 15.82 48.03 30.00 48.38
+H.1 Combination with Quantization
 Table 13: L-Eval scores of FullAttn, FullAttn-4bits, Locret and Locret-4bits. (Detailed)
@@
 The results in Table 13 shows that the degradation caused by quantization is not significantly high, showing that Locret exhibits good robustness on data representation and it is friendly to quantization.
-C.2 Combination with LoCoCo
+H.2 Combination with LoCoCo
 Table 14: L-Eval scores of LoCoCo, Locret and the combination LoCoCo+Locret. (Detailed)
@@
 In all methods, we set the local length to 0, following the original setting.
-Appendix D Training Robustness
+Appendix I Training Robustness
+Figure 6: L-Eval scores with different intermediate size of retaining head $d_{\mathbf{R}}$.
+Table 15: L-Eval scores of Locret trained on various datasets.
+Dataset LongAlpaca LongAlign Anti-Haystack
+L-Eval 21.33 22.00 20.72
 Locret demonstrates high robustness to the training settings, suggesting that there is no need for careful tuning of training hyperparameters or meticulous selection of datasets.
 Here, we ablate the intermediate size of the retaining heads $d_{\mathbf{R}}$ and train the retaining head $\mathbf{R}$s on various long-context tuning datasets to demonstrate the stability of results across different training settings.
-D.1 Intermediate Size of the retaining head
+I.1 Intermediate Size of the retaining head
 We align all the training settings as described in Section 4.1 and only change the intermediate size of retaining heads $d_{\mathbf{R}}\in\{256,512,1024,2048,4096\}$ with the backbone model Phi-3-mini-128K.
@@
 However, the performance variance is negligible compared to the increase of parameter size, thus we choose to maintain the intermediate size in a small scope to take balance of performance and efficiency.
-D.2 Training Data Insensitivity
-We also consider the sensitivity of the training data, which leads us to ablate the training dataset by training on LongAlign (Bai et al., 2024) and Anti-Haystack (Pan, 2024), comparing these results with those from LongAlpaca (Chen et al., 2024) in the original training setting.
+I.2 Training Data Insensitivity
+We also consider the sensitivity of the training data, which leads us to ablate the training dataset by training on LongAlign (Bai et al., 2024a) and Anti-Haystack (Pan, 2024), comparing these results with those from LongAlpaca (Chen et al., 2024) in the original training setting.
 We also align other settings to the original setting and choose the backbone model to be Phi-3-mini-128K.
@@
 Locret can obtain competitive performance without delicately selecting the training data.
-Appendix E Discontinuous Context and Stablizers
+Appendix J Extremely Long Context Evaluation
+[…]
+Table 18: Inference speed with Retaining Heads.
+Phi-3-mini-128K on 10M context
+Dataset R.PassKey_10M
+Locret 100.00
+Appendix K Compressing Multi-turn Conversations
+Compared to query-aware eviction methods, such as SnapKV (Li et al., 2024a), Locret is a more suitable solution for multi-turn conversation scenarios.
+This is because the evaluation of cache importance in Locret is based on the cache itself, rather than being dependent on the subsequent query.
+To demonstrate this, we evaluate Locret on the Rock-Paper-Scissors benchmark introduced in SirLLM (Yao et al., 2024).
+Since SirLLM is specifically designed for such scenarios, we use it as our baseline in this benchmark.
+Results in Table 19 show that Locret is also effective in multi-turn conversation contexts.
+The hyperparameters are aligned with those used in SirLLM, with the cache budget set to 1024, and no stabilizers are retained, as SirLLM does not retain local tokens in this benchmark.
+We perform 2000 turns as same as the original SirLLM settings.
+The results are presented below.
+Table 19: Rock-Paper-Scissors scores of Locret and SirLLM.
+Phi-3-mini-128K on Rock-Paper-Scissors
+Preference Rock Paper Scissors Avg.
+win tie lose win tie lose win tie lose $\uparrow$ $\downarrow$
+SirLLM 40.00 31.75 28.25 27.5 36.55 35.96 29.35 25.15 45.50 32.28 36.57
+Locret 18.95 50.00 31.05 30.35 19.45 50.20 52.05 27.25 20.70 33.78 33.98
+Appendix L Discontinuous Context and Stablizers
 Evicting cache units results in context discontinuity, which causes unstable CIS prediction and inaccurate calculation of later tokens.
@@
 This explains the reason for failure when the stabilizers are short or absent: context discontinuity leads to instability in the prediction of CIS, resulting in errors during cache eviction and amplifying errors in the hidden states.
-Appendix F Causal Importance Score Simulates a Cache Problem
+Appendix M Retaining Heads Do not Slow Down Inference
+We evaluate the model’s forward throughput under varying context lengths, both with and without retaining heads.
+The results are summarized below in Table 20.
+“$\mathbf{R}$” represents the retaining heads, and the throughput is reported in tokens per second (tok/s) in the format “Avg.
+/ Std.”
+Table 20: Inference speed with Retaining Heads.
+Context Length 1024 2048 3072 4096
+w/o $\mathbf{R}$ Speed 18674 / 443 19743 / 464 19982 / 402 20304 / 187
+w/ $\mathbf{R}$ Speed 17118 / 1117 18503 / 546 19054 / 283 19153 / 174
+From the results, no significant latency increase is observed when using retaining heads.
+The numerical differences are attributed to systematic variations rather than additional overhead introduced by retaining heads during inference.
+Appendix N Causal Importance Score Simulates a Cache Problem
 In this section, we show that assigning each cache unit a CIS and calculate each cache units with top-$b$ cache units simulates a cache problem, i.e.
@@
 Thus, Locret mathmatically equals to top-$b$ sparse attention.
-Definition F.1.
+Definition N.1.
 (Causal Calculation) Given a sequence of objects $c_{1},c_{2},\cdots,c_{n}$, if
@@
 […]
-Definition F.2.
+Definition N.2.
 (Causal Importance Score) Given a causal calculation $f$ and $c_{1},c_{2},\cdots,c_{n}$ is the generated sequence of $f$. $s_{1},s_{2},\cdots,s_{n}\in\mathbb{R}$ is a sequence of numbers. If
@@
 then $\{s_{i}\}$ is a CIS of sequence $\{c_{i}\}$. $h$ is a causal importance scoring function.
-Definition F.3.
+Definition N.3.
 (Cache Problem) Given a causal calculation $f=g\circ Sel$, its generated sequence $\{c_{i}\}$ and a positive number $b\in\mathbb{Z}_{+}$, if $f$ satisfies the following two condion, then $(f,b,\{c_{i}\})$ is a cache problem with budget $b$.
@@
   • $\forall 1\leq m_{1}<m_{2}\leq n,~{}Sel(c_{1},\cdots,c_{m_{2}})\backslash Sel(c_{1},\cdots,c_{m_{1}})\subseteq\{c_{m_{1}+1},\cdots,c_{m_{2}}\}.$
-Theorem F.1.
+Theorem N.4.
 (Calculating cache units with Top-$b$ CIS is a cache problem.) Given a causal calculation $f=g\circ Sel$, and its generated sequence $\{c_{i}\}$, a CIS $s_{i}=h(c_{i})$ and a positive number $b\in\mathbb{Z}_{+}$, if the selection function $Sel$ satisfies the following condition,
@@
 Thus, calculating cache units with Top-$b$ CIS is a cache problem.
-Appendix G Retained Patterns of Locret
+Appendix O Retained Patterns of Locret
 […]
@@
 In R.Number, we are able to observe a strong signal between token 10000 and 15000, which is the position of the inserted number string, indicating that Locret can identify the potentially answer-related parts by giving high predicted values of CIS.
-In M.Find, we can observe the StreamingLLM (Xiao et al., 2024b) pattern, where the tokens at the beginning of the sequence are always important.
+In M.Find, we can observe the StreamingLLM (Xiao et al., 2024c) pattern, where the tokens at the beginning of the sequence are always important.
 This is also mentioned as the $\Lambda$-pattern in MInference.
@@
 This pattern is also approached by MInference (Jiang et al., 2024a) by the pattern “vertical-and-slash”.
-In E.MC, H2(Zhang et al., 2024e) and ScissorHands (Liu et al., 2024a) pattern can be observed, following the assumption that if a token is activated at some point, it will continue to be activated in the consequencing process.
+In E.MC, H2O (Zhang et al., 2024d) and ScissorHands (Liu et al., 2024a) pattern can be observed, following the assumption that if a token is activated at some point, it will continue to be activated in the consequencing process.
 Noticing that the vertical lines always come in groups, which is the fundament of InfLLM (Xiao et al., 2024a) retrieving blocks to calculate.
@@
 Figure 8: Layer patterns of R.Number
-Appendix H The Training Process of Locret
+Appendix P The Loss Dynamics of Locret
 Here, we present changing trend of loss and accuracy during training in Figure 9.